### v1

# with Language Model Diagnostic Assistance

Abstract
In the emergency department (ED), patients undergo triage and multiple laboratory tests before diagnosis. This process is time-consuming, and causes ED crowding which significantly impacts patient mortality, medical errors, staff burnout, etc. This work proposes (time) cost-effective diagnostic assistance that explores the potential of artificial intelligence (AI) systems in assisting ED clinicians to make time-efficient and accurate diagnoses. Using publicly available patient data, we collaborate with ED clinicians to curate MIMIC-ED-Assist, a benchmark that measures the ability of AI systems in suggesting laboratory tests that minimize ED wait times, while correctly predicting critical outcomes such as death. We develop ED-Copilot which sequentially suggests patient-specific laboratory tests and makes diagnostic predictions. ED-Copilot uses a pre-trained bio-medical language model to encode patient information and reinforcement learning to minimize ED wait time and maximize prediction accuracy of critical outcomes. On MIMIC-ED-Assist, ED-Copilot improves prediction accuracy over baselines while halving average wait time from four hours to two hours. Ablation studies demonstrate the importance of model scale and use of a bio-medical language model. Further analyses reveal the necessity of personalized laboratory test suggestions for diagnosing patients with severe cases, as well as the potential of ED-Copilot in providing ED clinicians with informative laboratory test recommendations. Our code is available at https://github.com/cxcscmu/ED-Copilot.
AI for Health Care; Language Models for Emergency Department
1 Introduction
Emergency Department (ED) crowding represents a critical challenge in healthcare, significantly impacting morbidity, mortality, medical error, staff burnout, and incurring excessive costs (Sartini et al., 2022). Despite the documented effects of ED crowding, this issue remains underappreciated and inadequately addressed in healthcare systems. As a sentinel indicator of health system functioning, the ED is vital for providing timely access to care for patients who are seriously ill or injured (Savioli et al., 2022).
One key area to address ED crowding, as identified by the American College of Emergency Physicians, is to enhance throughput—the efficacy and efficiency of care delivery in the ED (Jarvis, 2016; DeAnda, 2018). A crucial factor affecting the throughput is long ED wait-times due to the laboratory testing process (Li et al., 2015). Patients frequently experience lengthy waits before their tests are ordered, followed by additional hours waiting for the test results. Clinicians can only make a diagnosis and decide on next steps for treatment after recieving these results. Furthermore, studies have indicated that approximately 40 to 60% of laboratory tests conducted in the ED are unnecessary (Miyakis et al., 2006), resulting in increased wait-times and reduced throuhput.
This paper proposes an artificial intelligence (AI) “Co-Pilot” system intended to offer (time) cost-effective diagnostic assistance in the ED. The system minimizes ED length of stay (LOS) – a proxy for wait-times – by automatically suggesting laboratory tests after patient triage. Minimizing wait-time via ED LOS serves to improve throughput, and hence crowding. Additionally, the system identifies patients at high risk of deterioration who require rapid intervention. By selecting informative laboratory tests, the system streamlines the diagnostic process, reducing time-cost (i.e., LOS) while ensuring excellent outcomes in terms of identifying high-risk patients.
To enable the machine learning (ML) community to study (time) cost-effective diagnostic assistance, we collaborate with ED clinicians and utilize publicly available patient records in MIMIC-IV (Johnson et al., 2023b) and other ED benchmarks (Xie et al., 2022) derived from MIMIC-IV, to develop MIMIC-ED-Assist, a benchmark that tests the ability of AI systems to help ED practitioners navigate the diagnostic space. Specifically, our ED clinicians identified two prediction targets: critical outcome, which includes patient death and ICU transfer (Levin et al., 2018), and lengthened ED stay, defined as ED LOS exceeding $24$ hours. Both reflect patient risk. Accurately identifying patients with these outcomes reduces time-cost by allowing clinicians to intervene in a timely fashion, and efficiently allocate resources. Furthermore, to mirror real-world ED practices, our clinicians grouped individual laboratory tests into commonly performed ED laboratory groups, e.g. complete blood count (CBC). MIMIC-ED-Assist then tests AI systems on their ability to recommend the most informative groups to make accurate diagnostic suggestions while minimizing the total time required to perform these tests, hence reducing LOS.
With MIMIC-ED-Assist, we develop ED-Copilot, which learns to suggest a series of laboratory groups, with minimum total time cost, and flags patients at high risk of critical outcome or lengthened ED stay. ED-Copilot first linearizes (i.e., converts to text) patient information, including demographic, triage, and laboratory test results into a text sequence. It then fine-tunes a bio-medical pre-trained language model BioGPT (Luo et al., 2022) to suggest future groups and predict the two outcomes. Next, we use a reinforcement learning (RL) framework introduced by Yu et al. (2023) to teach BioGPT to dynamically recommend the subsequent, most informative laboratory group based on prior lab and triage information. ED-Copilot effectively flags high-risk patients while minimizing the number of laboratory groups to reduce ED LOS, our proxy for wait-time.
Experiments on MIMIC-ED-Assist show that ED-Copilot is not only more accurate than state-of-the-art tree models trained on laboratory test results for both outcomes, but does so while using significantly fewer tests. ED-Copilot halves time cost of laboratory testing from four hours to two hours. Minimizing number of lab tests also has the benefit of reducing financial cost. We also perform ablation studies to confirm the benefits of our feature linearization technique and the bio-medical pre-trained language model backbone.
Our analyses confirmed the necessity of a personalized cost-effective diagnostic assistant. Non-personalized cost-effective methods minimize time-cost by selecting laboratory groups that maximize overall predictability. However, patients present differently, and require personalized care, particularly if they are at higher risk of critical outcome. By personalizing lab group suggestions, ED-Copilot provides more accurate diagnostic assistance for patients with unique needs. Furthermore, we show that ED-Copilot makes reasonable lab group suggestions despite results only available for tests the patient actually receives, due to the limitation of MIMIC-ED-Assist being a retrospective offline benchmark.
The main contributions of this paper are as follows.
  1. 1.
    We work with ED clinicians to develop MIMIC-ED-Assist, a publicly accessible benchmark designed to advance research in ED diagnostic assistance.
  2. 2.
    We develop ED-Copilot, an RL-trained language model that enhances ED workflow by suggesting informative laboratory groups and flagging high-risk patients using a pre-trained language model backbone BioGPT.
  3. 3.
    A myriad of experiments and analyses demonstrate that ED-Copilot improves accuracy over state-of-the-art classifier while reducing ED wait-time by 50%.
2 Related Work
Healthcare benchmarks. Researchers have spent considerable effort in converting raw electronic health records (EHRs) into large-scale open-source datasets to ensure easy access to high-quality medical data. A notable example is the Medical Information Mart for Intensive Care (MIMIC) database (Johnson et al., 2023b) which provides patient information such as measurements, laboratory orders, and treatments, ranging from the ED to inpatient care, including the intensive care unit (ICU). MIMIC has led to the development of a range of related prediction benchmarks and models (Harutyunyan et al., 2019; Purushotham et al., 2018; Wang et al., 2020) focused on the ICU. Xie et al. (2022) took a step towards filling this gap by using the MIMIC-IV-ED (Johnson et al., 2023a) database to build a ED-focused benchmark. Their dataset includes ED triage information, and various clinical outcomes such as hospitalization that interest clinicians, and impact ED LOS.
AI models for healthcare. Alongside developing benchmarks, the ML community has also tried to build models to predict various clinical outcomes. Traditional ML methods (e.g., random forests, gradient boosting, and their variants (Breiman, 2001; Chen & Guestrin, 2016; Agarwal et al., 2022, 2023)), along with deep learning (DL) have been used to predict pneumonia (Kang et al., 2020), and septic shock in the ICU (Wardi et al., 2021). Other works use interpretable models to provide diagnostic assistance in the ED such as identifying traumatic brain injury (Tan et al., 2022; Kornblith et al., 2022).
A line of work closely related to our paper focuses not only on prediction but also on AI for cost-effective medicine. For instance, Bejnordi et al. (2017) showed DL leads to faster analysis of pathology laboratory results; Komorowski et al. (2018) proposed an “AI clinician” to learn optimal dosing strategies for treating sepsis; Yu et al. (2023) focused on minimizing financial costs associated with laboratory testing while maximizing prediction accuracy. Specifically, Yu et al. (2023) used RL to sequentially select laboratory groups based on a patient’s observed test results to optimize this (financial) cost-accuracy trade-off. They validated the accuracy and cost-effectiveness of their approach on multiple clinical tasks such as diagnosing kidney injury. Researchers have also begun to explore the use of large language models (LLMs) in medical applications. LLMs have been used to extract clinical concepts (Luo et al., 2022; Yang et al., 2022, 2021), and facilitate medical question answering (Yagnik et al., 2024; Singhal et al., 2023). However, the integration of LLMs into ED workflows, particularly in addressing ED crowding, remains unexplored.
3 MIMIC-ED-Assist Benchmark
We discuss how we curate MIMIC-ED-Assist in collaboration with clinicians, a benchmark with comprehensive laboratory information that allows for the development of AI systems that can tackle long ED wait-times via minimizing laboratory testing time-cost.
Task description. We focus on the following two prediction tasks. 1) Flagging patients at high risk of critical outcome (i.e., death and ICU transfer), and lengthened ED. 2) Providing time cost-effective diagnostic assistance by recommending the next most medically informative laboratory group while simultaneously minimizing the time-cost of these groups. As discussed in detail below, both tasks help clinicians optimize their workflow, and serve to achieve the larger goal of reducing ED LOS.
Data pre-processing. We filter out non-hospitalized ED patients from MIMIC-IV-ED, as laboratory results are only available for admitted patients. We only focus on adults and remove patients younger than $18$ years old. Further, we exclude patients with missing triage information. This step is necessary since clinicians order laboratory groups depending on triage information. We simplify our task by removing patients who receive the same test multiple times, approximately $1.5\%$ of all patients.
Clinically relevant outcomes. We construct two prediction outcomes that help optimize ED workflow and reduce ED LOS. (1) Critical condition, which refers to death during hospitalization or transfer to an ICU within 12 hours. (2) Lengthened ED stay, indicating if ED LOS exceeding $24$ hours. These outcomes were chosen as prediction targets in close collaboration with ED clinicians for the following reasons. If a patient is identified to be at high-risk of a critical outcome, clinicians will prioritize treatment and resources for this patient. Lengthened ED stay is typically correlated with the complexity of a patient’s case. Flagging patients at high risk of lengthened ED stay can enable timely intervention, and reduce ED LOS. While these two outcomes are correlated, they cover different aspects of patient care. For example, patients at high risk of critical outcomes often should be hospitalized quickly, while patients with complications do not necessarily have severe cases. As such, healthcare providers often require different diagnostics and resource management for each of these two tasks.
Triage feature selection. We select a number of triage features in close collaboration with ED clinicians. Triage features are available for every patient before any laboratory tests are ordered. Specifically, our ED clinicians chose $9$ triage variables that are available at the beginning of the patient encounter. These variables included patient demographics, medical history, vital signs, and chief complaints from triage (i.e., natural language description of symptoms).
Laboratory test selection. We only include laboratory tests performed in the ED. For simplicity, we exclude laboratory tests received by less than $5\%$ of patients and leave further examination of rare laboratory tests in future research. This process results in a total of $68$ available laboratory tests. While there are $68$ laboratory tests, ED clinicians rarely order individual tests for a patient. Rather, they order groups or groups of tests based on a patient’s signs, symptoms, and risk factors. For example, a clinician might order a basic metabolic group or a complete blood count. To reflect this clinical practice, we work with ED clinicians to categorize these 68 tests into 12 distinct groups, representing common groupings in clinical settings. See Appendix A for all laboratory tests, and their assigned groupings. On average each patient has $4.7$ laboratory groups. Consequently, our dataset contains numerous missing values, each representing a laboratory test not administered to a specific patient.
ED LOS. Our benchmark records when each group is ordered, and assigns its average completion time as its ’time-cost’. ED LOS depends on the time-costs of the administered groups and can be modeled in different ways. For example, sequential tests result in ED LOS being the sum of time costs, whereas parallel tests imply ED LOS is equal to the group with the largest time-cost. MIMIC-ED-Assist does not specify how to approximate ED LOS from these time-costs, and instead provides this flexibility to researchers and practitioners.
Table 1: Statistics of MIMIC-ED-Assist. It includes information from patient triage and lab tests. Lab tests are grouped into 12 groups by ED clinicians based on how they are commonly ordered. An ED visit has a critical outcome if the patient is transferred to ICU or there is an inpatient mortality. ED stay is lengthened if the length of stay (LOS) exceeds 24 hours.
Variable/Label Count
##\## of ED visits 32356
##\## of patients 25714
##\## of triage variables 9
##\## of lab variables 67
##\## of lab groups 12
Avg ##\## of lab groups per patient 4.7
##\## of Inpatient mortality 467 (1.44%)
##\## of ICU transfer in 12h 2894 (8.94%)
##\## of Critical outcome 3129 (9.67%)
##\## of ED LOS $>$ 24h 2232 (6.90%)
Data availability. Our pipeline to curate MIMIC-ED-Assist from the MIMIC-IV datasets will be made available. MIMIC-IV is publicly available, and then one can use our pipeline to create MIMIC-ED-Assist.
Limitations. There are several limitations of MIMIC-ED-Assist, mainly due to data availability challenges in public patient data (1) MIMIC-ED-Assist is derived from MIMIC-IV which only has laboratory results for hospitalized ED patients. Thus, our dataset is not reflective of the entire population that visits the ED, but instead is biased towards those with more severe issues (hospitalized). (2) As an offline benchmark derived from past patient records, all data is retrospective. When training models to recommend future laboratory groups, one can only use laboratory tests patients actually received, otherwise no testing results are available. This leads to common challenges of offline benchmarks and may result in models learning sub-optimal laboratory group assignment, and unobserved confounding. (3) MIMIC-IV is collected at the Beth Israel Deaconess Medical Center, and may not reflect the distribution of other healthcare systems. Caution is needed when generalizing models and insights derived from MIMIC-IV to other places.
4 ED-Copilot for Diagnostic Assistance
With MIMIC-ED-Assist, we develop ED-Copilot, a method to provide (time)-cost-effective diagnostic assistance for ordering informative laboratory groups to flag patients at high risk for critical outcome and lengthened ED stay. We first provide a high-level description of ED-Copilot, followed by how it is trained. Training ED-Copilot involves a supervised fine-tuning stage for adapting the pre-trained language model (PLM) (e.g. BioGPT) backbone to our task, followed by an RL stage that selects laboratory groups to reduce time-cost. At the end of this section, we discuss how ED-Copilot conducts inference.
Figure 1: Overview of ED-Copilot training flow on one ED visit.
4.1 Problem Formulation
Task. The task for ED-Copilot is to assist doctors by recommending the next laboratory group to flag patients with potential critical outcome or lengthened ED stay, based on patient triage and laboratory information.
Let a patient have one triage group $x_{0}$ and $n$ laboratory groups $[x_{i}]_{i=1}^{n}$with their respective associated results $[r_{i}]_{i=0}^{n}$ in the order of $[x_{0},r_{0},x_{1},r_{1},x_{2},r_{2},....,x_{n},r_{n}]$, along with a diagnostic outcome $y$. ED-Copilot first linearizes this patient’s triage and laboratory into a text sequence, then passes it into a language model backbone to suggest the next lab group or predict the outcome.
[…]
Language Model Backbone. ED-Copilot uses a PLM $G_{\theta}$ to take the input sequence, append a [EOS] token after each group, and uses classification head upon the hidden representations of $G_{\theta}$ to select next laboratory group or predict the outcome:
$,\displaystyle[x_{0},r_{0},\texttt{[EOS]}_{0},$$...,x_{n},r_{n},\texttt{[EOS]}_{n},y]\xrightarrow{G_{\theta}}\textit{{H}},$ (1)
$p_{\phi}(x_{i}|\textit{{h}}_{<i})$ $=\text{MLP}_{\phi}(\textit{{h}}_{i-1}),$ (2)
$p_{\psi}(y|\textit{{h}}_{\leq n})$ $=\text{MLP}_{\psi}(\textit{{h}}_{n}).$ (3)
Here $G_{\theta}$ generates hidden representations for each token, on which we use a multi-layer perceptron (MLP) $p_{\phi}(x_{i}|\textbf{{h}}_{<i})$ to predict the next laboratory group from the laboratory label space and another MLP $p_{\psi}(y|\textbf{{h}}_{\leq n})$ to predict the diagnostic outcome $y$, conditioned on the hidden representation h of the previous [EOS] position. Note that we only predict laboratory groups, not their associated laboratory tests’ results, which should be filled by running the actual lab testing.
4.2 Supervised Fine-tuning
The first stage of training uses standard language model fine-tuning techniques, where we train the PLM to suggest the next laboratory group to order and eventually to predict target outcomes. To predict the next laboratory group, we use a standard auto-regressive loss function:
$\mathcal{L}_{\text{lab}}$ $=-\frac{1}{n}\sum_{i=1}^{n}\log p_{\phi}(x_{i}|\textit{{h}}_{<i}).$ (4)
For outcome prediction, we use the following loss
$\mathcal{L}_{y}$ $=-\log p_{\psi}(y|\textit{{h}}_{\leq n}).$ (5)
Then, we fine-tune the PLM by:
$\min_{\theta,\psi,\phi}(\mathcal{L}_{\text{lab}}+\mathcal{L}_{y}),$ (6)
where $\theta,\psi,\phi$ are parameters in the PLM $G_{\theta}$ and two MLPs $p_{\psi},p_{\phi}$. ED clinicians can use the fine-tuned PLM to suggest a sequence of laboratory groups and predict outcomes.
4.3 Reinforcement Learning
Choosing informative laboratory groups is a key factor in reducing laboratory testing, and hence improving wait-time, and reducing ED LOS. In this section, we employ RL to introduce the notion of time-cost effectiveness to the supervised fine-tuned PLM, to select laboratory groups that maximize F1-score while minimizing time-cost.
Markov Decision Process. The RL process can be viewed as a Markov Decision Process (MDP), represented by $\{\mathcal{S},\mathcal{A},P,R\}$. Let $\mathcal{S}$ denote the state space, and state $s\in\mathcal{S}$ corresponds to a patient’s observed triage information, and laboratory group results. Specifically, for a patient with observed information $[x_{0},r_{0},x_{1},r_{1},x_{2},r_{2},....,x_{n},r_{n}]$, let $s_{\leq i}=[x_{0},r_{0},x_{1},r_{1},x_{2},r_{2},....,x_{i},r_{i}]$ for $0\leq i\leq n$.
Action Space. The action space $\mathcal{A}=\{x_{1},x_{2},....,x_{K}\}\cup\{y^{+},y^{-}\}$ include $K$ number of groups, with action $a\in\{x_{1},x_{2},....,x_{K}\}$ corresponding to ordering a laboratory test group, with a time-cost $c(a)$. Action $a\in\{y^{+},y^{-}\}$ means predicting the outcome and terminating the process.
Policy. The policy $\pi_{\eta}:\mathcal{S}\rightarrow\mathcal{A}$ maps from states to actions, and is parameterized by weights $\eta$. Specifically, $\pi_{\eta}(a|\textbf{{h}}_{<i})$ outputs probability of an action $a$ using the hidden representation $\textbf{{h}}_{<i}$ corresponding to [EOS] token for state $s_{i}$.
RL Training. We train the policy $\pi_{\eta}$ to follow two objectives: maximize F1-score and minimize time-cost. We measure the time cost of policy $\pi_{\eta}$, as follows:
$\text{Cost}(\pi_{\eta})=\mathbb{E}_{\pi_{\eta}}[\sum_{t\geq 0}\sum_{j\in[K]}c(j)\cdot\mathbf{1}\{a_{t}=x_{j}\}],$ (7)
where $\mathbb{E}_{\pi_{\eta}}$ is the expectation under policy $\pi_{\eta}$. Using this definition of cost, let $\pi^{*}_{\eta}(\alpha,\beta)$ represent the policy that maximizes F1-score while minimizing time-cost for hyper-parameters $\alpha,\beta$ to be defined. Then, $\pi_{\alpha,\beta}^{*}$ is equivalent to the solution of the following program:
$\pi^{*}_{\eta}(\alpha,\beta)=\text{argmax}_{\pi_{\eta}}\{\text{TN}(\pi_{\eta})+\alpha\text{TP}(\pi_{\eta})+\beta\text{Cost}(\pi_{\eta})\}.$ (8)
TN$(\pi_{\eta})$ and TP$(\pi_{\eta})$ are the numbers of true negatives and true positives, under policy $\pi_{\eta}$. Hyper-parameters $\alpha,\beta$ control the trade-off between F1-score and time-cost. We train the MLP via proximal policy optimization (Schulman et al., 2017).
Measuring time-cost. We measure time-cost via the total time taken to run all laboratory groups. In the ED, laboratory groups are often ordered both in parallel, and sequentially since the decision on new tests to order depends on previous tests. Additionally, the number of tests an ED clinician orders also depends on other factors outside of the patient’s health record, such as insurance policy. As the first step towards reducing ED LOS with AI, this work uses the sum of test time as the optimization target, which is an approximation of the influence of test costs on total ED LOS. Better ways to model ED LOS are future research for both the AI and healthcare communities.
Table 2: Overall performances of ED-Copilot and baselines in predicting Critical Outcome (transfer to ICU or mortality) and Lengthened ED Stay (spending more than 24 hours in emergency room), and the Average Time Cost (in minutes) to make a prediction. Sensitivity and specificity are the true positive and true negative rates.
Critical Outcome Lengthened ED Stay
Model F1 AUC Sensitivity Specificity Avg. Time Cost F1 AUC Sensitivity Specificity Avg. Time Cost
Random Forest 0.377 0.807 0.754 0.748 265 Minutes 0.206 0.698 0.693 0.616 265 Minutes
XGBoost 0.379 0.807 0.731 0.744 265 Minutes 0.212 0.679 0.619 0.661 265 Minutes
LightGBM 0.394 0.813 0.725 0.769 265 Minutes 0.217 0.705 0.706 0.605 265 Minutes
3-layer DNN 0.339 0.743 0.676 0.683 265 Minutes 0.194 0.637 0.649 0.593 265 Minutes
SM-DDPO 0.353 0.780 0.685 0.763 182 Minutes 0.183 0.619 0.472 0.739 177 Minutes
ED-Copilot 0.413 0.820 0.750 0.779 125 Minutes 0.252 0.707 0.725 0.606 154 Minutes
4.4 Inference
During inference, ED-Copilot assists ED clinicians to optimize their workflow by suggesting the next most informative laboratory group to order, and also by flagging patients at high risk of critical outcome and lengthened ED stay. Specifically, given triage information and previous test results, ED-Copilot recommends additional tests. The results for these additional tests can then be fed back to ED-Copilot to suggest additional laboratory groups or to flag one of two possible outcomes. Further, ED-Copilot can be used in more flexible ways depending on a clinician’s needs. For example, multiple tests can be ordered using multiple suggestions from ED-Copilot.
(a) Critical Outcome F1
(b) Critical Outcome AUC
(c) Lengthened ED Stay F1
(d) Lengthened ED Stay AUC
Figure 2: Prediction accuracy and average number of laboratory groups of ED-Copilot with different maximum allowed times to perform lab tests. Each point reflects ED-Copilot’s F1/AUC (y-axes) at different time thresholds.
5 Experimental Set-up
Dataset Split. We randomly split the dataset using $80\%$ for training, $10\%$ for validation, and $10\%$ for testing, while ensuring each split has the same class distribution. The validation set is used to tune hyper-parameters for all methods. During inference, the initial state of that patient is set to be demographic and triage information.
Evaluation Metrics. We use four evaluation metrics: F1-score, area under the receiver operating characteristic (AUC), sensitivity (true positive rate), and specificity (true negative rate), all standard in healthcare tasks prediction tasks (Harutyunyan et al., 2019; Xie et al., 2022).
Baselines. For prediction tasks only, ED-Copilot is compared to three tree-based methods: random forests (Breiman, 2001), XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017), all of which are known to achieve strong performance on tabular data. Additionally, we compare ED-Copilot to a $3$-layer dense deep neural network (DNN).
Note that since each patient is not observed under all lab groups, there are many missing values in our dataset. For these baselines, we consider a number of imputation methods: (1) mean imputation, (2) median imputation, and (3) zero-imputation (i.e., replacing missing values by 00), (4) using a dummy indicator to encode missing values. Results with the best imputation method is presented in the main manuscript, and the rest are in Appendix C.
We also compare ED-Copilot to another cost-effective baseline, SM-DDPO (Yu et al., 2023), which selects laboratory groups to minimize time cost while maximizing F1-score. SM-DDPO is a non-personalized method, and selects the same laboratory groups for all patients. All hyper-parameters are tuned on the validation set to maximize F1-score.
Time Cost. Each group of lab tests is assigned a time cost by observing time-stamps in the MIMIC-IV database. For both ED-Copilot, and all baselines, we measure the amount of time taken by averaging the time required for an algorithm to make a prediction across all patients. This reflects the impact of laboratory tests on ED LOS.
Implementation Details. BioGPT (Luo et al., 2022) is used as our backbone language model, which is a generative pre-trained transformer (347 million parameters) for biomedical text generation and mining. In the RL phase of training ED-Copilot, for a given patient, our action space is restricted to laboratory groups that the patient actually performed. A discussion when selecting unobserved laboratory groups can be found in Section 6.4.
Table 3: Ablation study on components of ED-Copilot, including linearization techniques, feature importance, and PLM backbones.
Group Variations Critical Outcome Lengthened ED Stay
F1 AUC Sensitivity Specificity Avg. Time Cost F1 AUC Sensitivity Specificity Avg. Time Cost
ED-Copilot (345M) 0.413 0.820 0.750 0.779 125 Minutes 0.252 0.707 0.725 0.606 154 Minutes
Linearization Raw Lab Test Name 0.397 0.777 0.768 0.677 134 Minutes 0.241 0.695 0.611 0.701 144 Minutes
Features w/o. Triage 0.277 0.704 0.679 0.649 0.145 0.593 0.532 0.606
w/o. CBC 0.385 0.803 0.692 0.777 0.224 0.686 0.696 0.596
w/o. Chemical 0.420 0.827 0.788 0.746 0.234 0.702 0.656 0.606
Backbone Pythia (70M) 0.290 0.698 0.574 0.702 166 Minutes 0.178 0.596 0.555 0.619 126 Minutes
GPT2-Medium (345M) 0.358 0.757 0.621 0.747 133 Minutes 0.166 0.539 0.498 0.584 96 Minutes
6 Evaluation Results
6.1 Prediction Accuracy and Time Cost
Overall Performance. The prediction performance and time costs of all methods are presented in Table 2. For critical outcome, ED-Copilot outperforms the next best baseline by 1.9%, 0.7%, 2.5%, and 1% for F1-score, AUC, sensitivity, and specificity, respectively. For lengthened ED stay, ED-Copilot outperforms the next best model by 3.5%, 0.2%, 1.9%, and 0.1% for the four metrics. ED-Copilot halves average time cost from roughly $4$ hours to $2$ hours.
In comparison to the other cost-effective baseline, SM-DDPO, ED-Copilot also has significantly better accuracy and lower time cost. Both methods utilize the same RL algorithm to minimize time-cost, but ED-Copilot benefits from the strong capability of PLM backbone (Sec 6.2) and the personalized assistance it enabled (Sec 6.3).
Prediction Accuracy at Different Time Costs. In Figure 2, we investigate how ED-Copilot’s prediction performance changes as we vary the maximum time allowed to perform laboratory tests. The prediction performance increases with the maximum allowed time and number of laboratory groups selected. ED-Copilot requires relatively few tests to approximate peak performance for lengthened ED stay, while prediction accuracy for critical outcome steadily increases with more laboratory tests allowed. These curves illustrate different trade-offs of diagnostic assistance for different outcomes and help ED clinicians better leverage when allocating resources.
(a) Sensitivity-Specificity $\alpha$
(b) F1-Cost $\beta$
Figure 3: Impact of Hyper-parameters on Sensitivity-Specificity ($\alpha$) and F1-Cost ($\beta$) trade-off when predicting critical outcome.
6.2 Ablation Studies
PLM Backbone. We compare BioGPT to GPT2 at the same size, and a smaller scale model Pythia (Biderman et al., 2023) that has 70 million parameters. BioGPT is based on GPT2 (Radford et al., 2019), but is further pre-trained on a bio-medical corpus. Table 3 shows both model scale and bio-medical pre-training lead to stronger performance.
Linearization Technique. As done in Hegselmann et al. (2023); Manikandan et al. (2023), we replace the true laboratory test name with a standard feature name while keeping the result for that laboratory group unchanged. That is, we linearize laboratory results as follows ( feature1 : value1 |normal-||| feature2 : value2 … ). Table 3 shows using standard feature names leads to better F1 and AUC score as compared to using the raw laboratory test names. One possible reason is that using raw laboratory test names biases ED-Copilot to pick names over informativeness for prediction.
Feature Importance. We examine the change in ED-Copilot performance when training with and without a given feature set from top three popular feature sets: triage information, complete blood count, and chemical group. Table 3 shows removing triage features leads to a significant drop in performance. Our ED collaborators confirmed this made medical sense since initial laboratory tests are based on triage information. Our results also show all laboratory groups are not equally important, with removing the CBC causing a larger drop in performance than excluding the chemical laboratory groups.
Figure 4: Fraction of patients performing lab groups and predicted by ED-Copilot. On average each patient performed 4.7 groups and cost-effective ED-Copilot suggested 2.4 groups.
Hyper-parameters. Hyper-parameters $(\alpha,\beta)$ in Equation 8 control the trade-off between prediction accuracy and time-cost in training. Increasing $\alpha$ trades off sensitivity over specificity, while increasing $\beta$ trades off F1-score over time-cost. Figure 2(a) shows increasing $\alpha$ leads to trade off sensitivity over specificity. Figure 2(b) show increasing $\beta$ causes ED-Copilot to trade off F1-score over time-cost. The hyper-parameters $(\alpha,\beta)$ provide a control for ED clinicians to change ED-Copilot’s behavior to fit their preferences.
6.3 Analysis on Personalized Diagnostic Assistance
Table 4: Critical Outcome prediction accuracy at different cohorts, by the category of lab groups they performed (Top: Performed at least one of the top two most frequent lab groups; Middle: Performed at least one of the middle six lab groups; Rare: Performed one or more rare tests in the last four lab groups). The total number of positive (critical)/negative cases and positive rate is shown in parentheses.
W. Top 2 Lab Groups (302/2823,9.6% ) W. Middle 6 Lab Groups (299/2603, 10.3%) W. Last 4 Lab Groups (141/817, 14.7%)
Model F1 Sensitivity Specificity F1 Sensitivity Specificity F1 Sensitivity Specificity
Random Forest (Top 2 groups) 0.330 0.735 0.752 0.335 0.746 0.750 0.405 0.730 0.739
XGBoost (Top 2 groups) 0.361 0.788 0.680 0.374 0.732 0.728 0.433 0.738 0.718
LightGBM (Top 2 groups) 0.401 0.788 0.705 0.409 0.806 0.690 0.462 0.738 0.742
SM-DDPO 0.364 0.760 0.715 0.373 0.762 0.724 0.435 0.732 0.734
ED-Copilot 0.414 0.701 0.788 0.431 0.731 0.774 0.461 0.767 0.720
Personalized treatment strategies are often required since patients have diverse needs and present differently. ED-Copilot learns a personalized representation for each patient and naturally supports personalized diagnostic assistance. On the other hand, previous cost-effective methods based on traditional ML algorithms (e.g., tree-based methods), optimize cost-effectiveness at the full population level, i.e., via feature selection/pruning, rather than personalized for each patient. However, this does not naturally support personalization, unlike ED-Copilot which suggests individualized future tests based on a patient’s triage and previous test results. The following experiments highlight the benefit of personalization via ED-Copilot .
First, we plot both the fraction of patients receiving each group as well as the fraction of patients predicted by ED-Copilot in Figure 4. In the ED, patients undergo a large variety of laboratory groups: after the two most common groups (Chem and CBC), more than half of the patients performed some other tests. ED-Copilot predicts fewer laboratory groups than the observed data since it is trying to minimize time-cost, yet it still predicts a wide variation of recommended tests depending on the patient’s condition. On the other hand, non-personalized cost-effective methods, often withdraw to picking the most frequent groups.
To further understand the benefit of personalization, we partition patients in the test set into three cohorts based on the rarity of the group they were administered; see Table 4 for a precise definition of these cohorts. The middle and rare cohorts have more positive cases, indicating higher severity in these cohorts. We train tree-based baselines using triage information and results from the two most frequent groups (Chem and CBC), and compare it to ED-Copilot. Results in Table 4 show that ED-Copilot and baselines perform comparably on patients in the top cohorts (i.e., patients that present typically). For the middle cohort, ED-Copilot greatly improves F1-score as well. Further, in the rare cohort which has the highest fraction of critical cases as compared to other groups (15% versus 10%), ED-Copilot achieves higher significantly higher sensitivity than other methods. Hence, ED-Copilot provides personalized assistance to provide more equitable care, particularly for high-risk patients.
(a) Accuracy
(b) Avg. # of Labs
Figure 5: Comparison of prediction performance on critical outcome of ED-Copilot when restricted or not restricted to tests patients performed in its suggestion at different time cost constraint.
6.4 Unrestricted Lab Group Suggestion
Since MIMIC-ED-Assist is an offline retrospective benchmark, we restrict ED-Copilot during training to only select laboratory groups that patients have received. However, this restriction may not lead to the optimal group suggestion. This experiment investigates ED-Copilot’s performance without this restriction.
Specifically, if ED-Copilot (unrestricted) selects a laboratory group not received by a patient, we impute those results by 00, but add to the time-cost. ED-Copilot (unrestricted) achieves an accuracy of $77\%$ when predicting the next laboratory group, i.e., the group actually received by a given patient, indicating that ED-Copilot (unrestricted) is reasonably good at fitting the clinician’s decision but there are still differences between model suggestions and historical data.
Next, we compare ED-Copilot (unrestricted) to ED-Copilot (restricted) in terms of F1-score for critical outcome when varying the maximum allowed time. Figure 5 show that for low maximum allowed time, ED-Copilot (unrestricted) performs worse than ED-Copilot since if it picks unobserved labs, the cost brings no new information added (no results available). As the number of allowed laboratory groups increases, ED-Copilot (unrestricted) matches the performance of ED-Copilot. Restriction to observed laboratory groups might be necessary for offline evaluation. When used in real scenarios, ED-Copilot would serve as a “Co-Pilot” and suggest laboratory groups from the entire set without restriction, assisting ED clinicians to optimize their workflows.
7 Conclusion
In this paper, we aim at reducing emergency department wait time through time-cost-effective diagnostic assistance, in collaboration with ED clinicians. Specifically, we curate MIMIC-ED-Assist, a benchmark that contains comprehensive laboratory information annotated with medically relevant groupings alongside key clinical outcomes that serve as useful approximations of ED wait-time. Using MIMIC-ED-Assist, we propose ED-Copilot, an AI system that provides time-cost-effective diagnostic assistance for ED clinicians, by recommending informative laboratory groups, and flagging patients at high risk of critical outcome and lengthened ED stay. ED-Copilot outperforms baselines by significantly improving prediction accuracy and reducing laboratory time costs. Further, its language model backbone allows personalized diagnostic assistance to better address the needs of patients who present atypically. We believe that this work serves as a stepping stone towards AI-driven solutions in the ED, and we hope that MIMIC-ED-Assist will spur interest in applying advancements in ML and language models to tackle this critical healthcare challenge.
8 Acknowledgments
We gratefully acknowledge partial support from NSF grants DMS-2209975, 2015341, NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI), the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and 814639, and NIH grant R01GM152718. We also acknowledge support by the Eunice Kennedy Shriver National Institute of Child Health and Human Development of the National Institutes of Health under Award Number K23HD110716. We also thank Yikuan Li from Northwestern University to share their preprocessed sepsis dataset.
Appendix A Triage and Lab Test Group
In Table 5, we include triage assessments and 12 lab groups with 68 tests and their estimated time cost.
Table 5: Triage and Lab Test Group
Triage Age Diastolic blood pressure
Gender Self-reported pain
Heart rate Emergency severity index acuity
Respiratory rate Chief complaint
Systolic blood pressure Oxygen saturation
Temperature Past commodities
Past ICU/ED/hospital visiting frequency
Complete Blood Count (CBC) (30 min) Hematocrit Neutrophils
White Blood Cells Red Cell Distribution Width (Standard Deviation)
Hemoglobin Absolute Lymphocyte Count
Red Blood Cells Absolute Basophil Count
Mean Corpuscular Volume Absolute Eosinophil Count
Mean Corpuscular Hemoglobin Absolute Monocyte Count
Mean Corpuscular Hemoglobin Concentration Absolute Neutrophil Count
Red Blood Cell Distribution Width Bands
Platelet Count Atypical Lymphocytes
Basophils Nucleated Red Cells
Eosinophils Monocytes
Lymphocytes
Chemistry (CHEM) (60 min) Urea Nitrogen Glucose (Chemistry)
Creatinine Potassium
Sodium Anion Gap
Chloride Calcium, Total
Bicarbonate
Coagulation (COAG)(48 min) Prothrombin Time Partial thromboplastin time
International Normalised Ratio
Urinalysis (UA) (40 min) PH (Urine) Protein
Specific Gravity Hyaline Casts
Red Blood Count (Urine) Ketone
White Blood Count (Urine) Urobilinogen
Epithelial Cells Glucose (Urine)
Lactate (4 min) Lactate
Liver Function (LFTs)(104 min) Alkaline Phosphatase Bilirubin, Total
Asparate Aminotransferase (AST) Albumin
Alanine Aminotransferase (ALT)
Lipase (100 min) Lipase
Electrolyte (LYTES)(89 min) Magnesium Phosphate
Cardiovascular (CARDIO) (122 min) NT-proBNP Troponin T
Blood Gas (12 min) Potassium, Whole Blood PO2
PH (Blood Gas) PCO2
Calculated Total CO2 Glucose (Blood Gas)
Base Excess Sodium, Whole Blood
Toxicology (TOX) (70 min) Ethanol
Inflammation (178 min) Creatine Kinase (CK) C-Reactive Protein
Appendix B Training details
We use a $3$-layer neural network for all MLPs in ED-Copilot. Due to large class imbalance, we use class weights when training the diagnostic predictor $p_{\psi}$. All experiments, training and hyper-parameter tuning are conducted on a NVIDIA RTX A6000 GPU. During the RL phase of training ED-Copilot, we restrict the action space to be on laboratory groups that were administered to patients. That is, we do not select laboratory groups that patients have not received. Performance when this restriction is removed is explored in Section 6.3. To ensure ensure sufficient policy sampling on experience replay buffer and obtain a large batch size, we freeze the weights of our PLM, to train the policy $\pi_{\eta}$, which is parameterized by $\eta$. We use the PPO algorithm (Schulman et al., 2017) to train $\pi_{\eta}$. The loss function employed is as the following for hyper-parameters $c_{1}$, and $c_{2}$:
$\mathcal{L}_{\text{rl}}(\eta;\eta^{old})$ $=\mathcal{L}_{\text{clip}}(\eta;\eta^{old})-c_{1}\mathcal{L}_{\text{value}}(\eta)+c_{2}\text{Entropy}[\pi_{\eta}],$
$\mathcal{L}_{\text{clip}}(\eta;\eta^{old})$ $={\hat{\mathbb{E}}}_{i}\left[\min\left(\frac{\pi_{\eta}(a_{i}|{\textbf{{h}}}_{<i})}{\pi_{\eta^{old}}(a_{i}|{\textbf{{h}}}_{<i})}\hat{A}_{i},\text{clip}\left(\frac{\pi_{\eta}(a_{i}|{\textbf{{h}}}_{<i})}{\pi_{\eta^{old}}(a_{i}|{\textbf{{h}}}_{<i})},1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right)\right],$
$\mathcal{L}_{\text{value}}(\eta)$ $=\hat{\mathbb{E}}_{i}\left[(V_{\eta}({\textbf{{h}}}_{<i})-V_{i})^{2}\right],$
Here ${\textbf{{h}}}_{<i}$ is the hidden representation from PLM of last [EOS] in this patient’s linearized sequence in state $s_{i}$. $\mathcal{L}_{\text{clip}}$ is clipped surrogate loss and $\hat{A}_{i}$ is estimated advantages which is regularized by value function $\mathcal{L}_{\text{value}}$. $\text{Entropy}[\pi_{\eta}]$ denotes an entropy term over the states, and $\hat{\mathbb{E}}_{i}$ is the empirical average over the collected dataset. We used a masked actor-critic network with package stable-baseline3 (Raffin et al., 2021). See algorithm 1 for a high-level description of this method.
Algorithm 1 Proximal Policy Optimization (PPO)
  for iteration $=0,1,\ldots$ do
     for actor $=1,2,\ldots,N_{\text{actor}}$ do
        Run policy $\pi_{\eta_{\text{old}}}$ in environment for $T$ timesteps and save all observations in experience buffer
        Compute estimated advantage $\hat{A}_{1},\ldots,\hat{A}_{i}$
     end for
     Optimize surrogate $\mathcal{L}_{\text{rl}}$ w.r.t $\eta$, with $D$ epochs and minibatch size $M\leq N_{\text{actor}}T$
     $\eta_{\text{old}}\leftarrow\eta$
  end for
We list the hyper-parameters in Table 6, including the supervised fine-tuning and reinforcement learning stage. In the RL stage, we use grid-search to tune $\alpha$ and $\beta$ to balance the trade-off between accuracy and cost. The search scope for $\alpha\in\{\frac{1}{8},\frac{1}{4},\frac{1}{2},1,2,4,8,15,16,32,64,256\}$ and $\beta\in\{\frac{1}{100},\frac{1}{50},\frac{1}{20},\frac{1}{10},1,10,100,1000\}$.
The parameters in PPO that are not specified are assigned the default values found in the Python package (Schulman et al., 2017).
Table 6: Hyperparameters Configurations for ED-Copilot
Hyperparameter Supervised Fine-tuning Reinforcement Learning
Learning Rate 1e-5
DNN Hidden size (3-layer) 1024
Optimizer AdamW
Adam $\epsilon$ 1e-8
Adam Betas ($\beta_{1}$, $\beta_{2}$) (0.9, 0.999)
Weight decay 0.01
Batch Size 32 128
Epochs 15 10
Max sequence length 768 656
Class Weight 10 -
Warmup percentage 0.1 -
Buffer Steps - 2048
Timesteps PPO trained per epoch - 20000
The penalty ratio between false positive and false negative $\alpha$ 15
The penalty ratio between false positive and lab cost $\beta$ $\frac{1}{100}$
Appendix C Baseline Details
In Table 7, we present tree models’ results with different imputation methods: (1) mean imputation, (2) median imputation, and (3) zero-imputation (i.e., replacing missing values by 0), (4) using a dummy indicator to encode missing values.
Table 7: Results with different imputation methods
Model Method Critical Outcome Lengthened ED Stay
F1 AUC Sensitivity Specificity F1 AUC Sensitivity Specificity
Random Forest Mean 0.377 0.807 0.754 0.748 0.198 0.701 0.736 0.578
Median 0.355 0.807 0.793 0.709 0.206 0.698 0.693 0.616
Zero 0.367 0.808 0.738 0.748 0.194 0.702 0.628 0.674
XGBoost Mean 0.328 0.768 0.715 0.697 0.212 0.678 0.593 0.683
Median 0.358 0.783 0.731 0.708 0.197 0.674 0.658 0.601
Zero 0.374 0.818 0.735 0.754 0.212 0.680 0.619 0.662
Dummy 0.379 0.807 0.731 0.744 0.190 0.656 0.680 0.557
LightGBM Mean 0.390 0.812 0.764 0.732 0.217 0.705 0.707 0.606
Median 0.394 0.813 0.725 0.769 0.190 0.681 0.688 0.599
Zero 0.393 0.814 0.735 0.774 0.209 0.702 0.697 0.623
Dummy 0.379 0.807 0.731 0.744 0.192 0.671 0.654 0.629

### v2

@@
 In the emergency department (ED), patients undergo triage and multiple laboratory tests before diagnosis.
-This process is time-consuming, and causes ED crowding which significantly impacts patient mortality, medical errors, staff burnout, etc.
-This work proposes (time) cost-effective diagnostic assistance that explores the potential of artificial intelligence (AI) systems in assisting ED clinicians to make time-efficient and accurate diagnoses.
-Using publicly available patient data, we collaborate with ED clinicians to curate MIMIC-ED-Assist, a benchmark that measures the ability of AI systems in suggesting laboratory tests that minimize ED wait times, while correctly predicting critical outcomes such as death.
-We develop ED-Copilot which sequentially suggests patient-specific laboratory tests and makes diagnostic predictions.
-ED-Copilot uses a pre-trained bio-medical language model to encode patient information and reinforcement learning to minimize ED wait time and maximize prediction accuracy of critical outcomes.
+This time-consuming process causes ED crowding which impacts patient mortality, medical errors, staff burnout, etc.
+This work proposes (time) cost-effective diagnostic assistance that leverages artificial intelligence systems to help ED clinicians make efficient and accurate diagnoses.
+In collaboration with ED clinicians, we use public patient data to curate MIMIC-ED-Assist, a benchmark for AI systems to suggest laboratory tests that minimize wait time while accurately predicting critical outcomes such as death.
+With MIMIC-ED-Assist, we develop ED-Copilot which sequentially suggests patient-specific laboratory tests and makes diagnostic predictions.
+ED-Copilot employs a pre-trained bio-medical language model to encode patient information and uses reinforcement learning to minimize ED wait time and maximize prediction accuracy.
 On MIMIC-ED-Assist, ED-Copilot improves prediction accuracy over baselines while halving average wait time from four hours to two hours.
-Ablation studies demonstrate the importance of model scale and use of a bio-medical language model.
-Further analyses reveal the necessity of personalized laboratory test suggestions for diagnosing patients with severe cases, as well as the potential of ED-Copilot in providing ED clinicians with informative laboratory test recommendations.
+ED-Copilot can also effectively personalize treatment recommendations based on patient severity, further highlighting its potential as a diagnostic assistant.
+Since MIMIC-ED-Assist is a retrospective benchmark, ED-Copilot is restricted to recommend only observed tests.
+We show ED-Copilot achieves competitive performance without this restriction as the maximum allowed time increases.
 Our code is available at https://github.com/cxcscmu/ED-Copilot.
@@
 Emergency Department (ED) crowding represents a critical challenge in healthcare, significantly impacting morbidity, mortality, medical error, staff burnout, and incurring excessive costs (Sartini et al., 2022).
-Despite the documented effects of ED crowding, this issue remains underappreciated and inadequately addressed in healthcare systems.
-As a sentinel indicator of health system functioning, the ED is vital for providing timely access to care for patients who are seriously ill or injured (Savioli et al., 2022).
+Despite the documented effects of ED crowding, this issue remains inadequately addressed in healthcare systems.
+An efficient and effective ED is vital for providing timely care to severely ill or injured patients (Savioli et al., 2022).
 One key area to address ED crowding, as identified by the American College of Emergency Physicians, is to enhance throughput—the efficacy and efficiency of care delivery in the ED (Jarvis, 2016; DeAnda, 2018).
-A crucial factor affecting the throughput is long ED wait-times due to the laboratory testing process (Li et al., 2015).
-Patients frequently experience lengthy waits before their tests are ordered, followed by additional hours waiting for the test results.
-Clinicians can only make a diagnosis and decide on next steps for treatment after recieving these results.
-Furthermore, studies have indicated that approximately 40 to 60% of laboratory tests conducted in the ED are unnecessary (Miyakis et al., 2006), resulting in increased wait-times and reduced throuhput.
-This paper proposes an artificial intelligence (AI) “Co-Pilot” system intended to offer (time) cost-effective diagnostic assistance in the ED.
-The system minimizes ED length of stay (LOS) – a proxy for wait-times – by automatically suggesting laboratory tests after patient triage.
-Minimizing wait-time via ED LOS serves to improve throughput, and hence crowding.
-Additionally, the system identifies patients at high risk of deterioration who require rapid intervention.
-By selecting informative laboratory tests, the system streamlines the diagnostic process, reducing time-cost (i.e., LOS) while ensuring excellent outcomes in terms of identifying high-risk patients.
-To enable the machine learning (ML) community to study (time) cost-effective diagnostic assistance, we collaborate with ED clinicians and utilize publicly available patient records in MIMIC-IV (Johnson et al., 2023b) and other ED benchmarks (Xie et al., 2022) derived from MIMIC-IV, to develop MIMIC-ED-Assist, a benchmark that tests the ability of AI systems to help ED practitioners navigate the diagnostic space.
-Specifically, our ED clinicians identified two prediction targets: critical outcome, which includes patient death and ICU transfer (Levin et al., 2018), and lengthened ED stay, defined as ED LOS exceeding $24$ hours.
-Both reflect patient risk.
-Accurately identifying patients with these outcomes reduces time-cost by allowing clinicians to intervene in a timely fashion, and efficiently allocate resources.
-Furthermore, to mirror real-world ED practices, our clinicians grouped individual laboratory tests into commonly performed ED laboratory groups, e.g.
-complete blood count (CBC).
-MIMIC-ED-Assist then tests AI systems on their ability to recommend the most informative groups to make accurate diagnostic suggestions while minimizing the total time required to perform these tests, hence reducing LOS.
-With MIMIC-ED-Assist, we develop ED-Copilot, which learns to suggest a series of laboratory groups, with minimum total time cost, and flags patients at high risk of critical outcome or lengthened ED stay.
+A crucial factor affecting throughput is the laboratory testing process, where patients often face lengthy waits for tests to be ordered and completed, delaying diagnoses and treatment decisions  (Li et al., 2015).
+Studies also show that 40 to 60% of ED laboratory tests are unnecessary (Miyakis et al., 2006), further exacerbating wait times.
+This paper proposes an artificial intelligence “Co-Pilot” system intended to offer (time) cost-effective diagnostic assistance in the ED.
+This system should aid diagnosis and minimize ED length of stay (LOS), i.e., wait times, by suggesting laboratory tests after patient triage.
+Further, it should help with resource management and planning by identifying severely ill patients who require rapid intervention.
+That is, by selecting informative tests, the system streamlines the diagnostic process, reducing LOS while improving outcomes, particularly for high-risk patients.
+To support the machine learning (ML) community in developing a time-cost-effective diagnostic assistant, we collaborate with ED clinicians to curate a benchmark, called MIMIC-ED-Assist, that is derived from MIMIC-IV (Johnson et al., 2023b) and related datasets (Xie et al., 2022).
+MIMIC-ED-Assist is designed to test the ability of AI systems to provide both accurate and time-cost saving laboratory recommendations.
+Our benchmark consists of two prediction targets identified by our clinical collaborators to reflect patient risk: critical outcomes which includes patient death and ICU transfer (Levin et al., 2018), and lengthened ED stay, defined as ED LOS exceeding $24$ hours.
+Accurately identifying patients at high risks of these outcomes reduces time-cost by allowing clinicians to perform timely interventions and efficiently allocate resources.
+MIMIC-ED-Assist mirrors real-world ED practices by grouping individual laboratory tests into commonly performed groups, e.g., complete blood count (CBC).
+MIMIC-ED-Assist then tests AI systems on their ability to recommend the most informative groups to make accurate diagnostic suggestions while minimizing the total time required to perform these tests, thereby reducing LOS.
+With MIMIC-ED-Assist, we propose ED-Copilot which suggests a series of laboratory groups to flag patients at high risks on our prediction targets while minimizing total time-cost.
 ED-Copilot first linearizes (i.e., converts to text) patient information, including demographic, triage, and laboratory test results into a text sequence.
-It then fine-tunes a bio-medical pre-trained language model BioGPT (Luo et al., 2022) to suggest future groups and predict the two outcomes.
-Next, we use a reinforcement learning (RL) framework introduced by Yu et al.
-(2023) to teach BioGPT to dynamically recommend the subsequent, most informative laboratory group based on prior lab and triage information.
-ED-Copilot effectively flags high-risk patients while minimizing the number of laboratory groups to reduce ED LOS, our proxy for wait-time.
-Experiments on MIMIC-ED-Assist show that ED-Copilot is not only more accurate than state-of-the-art tree models trained on laboratory test results for both outcomes, but does so while using significantly fewer tests.
-ED-Copilot halves time cost of laboratory testing from four hours to two hours.
-Minimizing number of lab tests also has the benefit of reducing financial cost.
+It then fine-tunes a bio-medical pre-trained language model BioGPT (Luo et al., 2022) to suggest future groups and predict our two defined targets.
+Next, we use a reinforcement learning (RL) framework (Yu et al., 2023) to teach BioGPT to dynamically recommend the subsequent, most informative laboratory group based on prior laboratory and triage information.
+Unlike baselines, ED-Copilot serves as a personalized diagnostic assistant since it uses past patient information to recommend future medically relevant laboratory groups.
+Experiments on MIMIC-ED-Assist show that ED-Copilot outperforms state-of-the-art tree models while halving time-costs of laboratory testing from four hours to two hours.
+Reducing the number of laboratory tests also has the benefit of reducing financial cost.
 We also perform ablation studies to confirm the benefits of our feature linearization technique and the bio-medical pre-trained language model backbone.
-Our analyses confirmed the necessity of a personalized cost-effective diagnostic assistant.
-Non-personalized cost-effective methods minimize time-cost by selecting laboratory groups that maximize overall predictability.
-However, patients present differently, and require personalized care, particularly if they are at higher risk of critical outcome.
-By personalizing lab group suggestions, ED-Copilot provides more accurate diagnostic assistance for patients with unique needs.
-Furthermore, we show that ED-Copilot makes reasonable lab group suggestions despite results only available for tests the patient actually receives, due to the limitation of MIMIC-ED-Assist being a retrospective offline benchmark.
-The main contributions of this paper are as follows.
-  1. 1.
-    We work with ED clinicians to develop MIMIC-ED-Assist, a publicly accessible benchmark designed to advance research in ED diagnostic assistance.
-  2. 2.
-    We develop ED-Copilot, an RL-trained language model that enhances ED workflow by suggesting informative laboratory groups and flagging high-risk patients using a pre-trained language model backbone BioGPT.
-  3. 3.
-    A myriad of experiments and analyses demonstrate that ED-Copilot improves accuracy over state-of-the-art classifier while reducing ED wait-time by 50%.
+Our ablation studies also investigate the effect of size of the language model backbone on prediction accuracy which indicates larger models can lead to further gain in performance.
+Our analyses also confirm the benefit of ED-Copilot’s personalized modeling approach.
+We show ED-Copilot can adapt its recommendations based on patient severity, thereby providing more accurate diagnostic suggestions for severely ill patients as compared to non-personalized baselines.
+Further, ED-Copilot achieves consistent performance across various subgroups such as age and sex.
+Lastly, since MIMIC-ED-Assist is a retrospective offline benchmark, we restrict ED-Copilot to only select laboratory tests a patient actually receives.
+We perform simulations without this restriction to approximate online performance, and show ED-Copilot is still able to make medically appropriate recommendations.
+The rest of this paper is organized as follows.
+In Section 2, we review related work.
+We discuss MIMIC-ED-Assist and ED-Copilot in Sections 3 and 4 respectively.
+Sections 5 and 6 discuss our experimental set-up and results.
 2 Related Work
@@
 A notable example is the Medical Information Mart for Intensive Care (MIMIC) database (Johnson et al., 2023b) which provides patient information such as measurements, laboratory orders, and treatments, ranging from the ED to inpatient care, including the intensive care unit (ICU).
-MIMIC has led to the development of a range of related prediction benchmarks and models (Harutyunyan et al., 2019; Purushotham et al., 2018; Wang et al., 2020) focused on the ICU.
+MIMIC has led to the development of a range of related prediction benchmarks and models (Purushotham et al., 2018; Harutyunyan et al., 2019; Wang et al., 2020) focused on the ICU.
 Xie et al.
@@
 AI Models for Healthcare.
-Alongside developing benchmarks, the ML community has also tried to build models to predict various clinical outcomes.
-Traditional ML methods (e.g., random forests, gradient boosting, and their variants (Breiman, 2001; Chen & Guestrin, 2016; Agarwal et al., 2022, 2023)), along with deep learning (DL) have been used to predict pneumonia (Kang et al., 2020), and septic shock in the ICU (Wardi et al., 2021).
-Other works use interpretable models to provide diagnostic assistance in the ED such as identifying traumatic brain injury (Tan et al., 2022; Kornblith et al., 2022).
-A line of work closely related to our paper focuses not only on prediction but also on AI for cost-effective medicine.
+There has been significant effort to apply ML to accurately predict clinical outcomes.
+Traditional methods (e.g., random forests, gradient boosting, and their variants (Breiman, 2001; Chen & Guestrin, 2016; Agarwal et al., 2022, 2023)), along with deep learning (DL) have been used to predict pneumonia (Kang et al., 2020), and septic shock in the ICU (Wardi et al., 2021).
+Other works use interpretable models to provide diagnostic assistance in the ED such as identifying traumatic brain injury (Kornblith et al., 2022; Tan et al., 2022).
+Another closely related line of work is AI for cost-effective medicine.
 For instance, Bejnordi et al.
-(2017) showed DL leads to faster analysis of pathology laboratory results; Komorowski et al.
+(2017) showed DL led to faster analysis of pathology laboratory results; Komorowski et al.
 (2018) proposed an “AI clinician” to learn optimal dosing strategies for treating sepsis; Yu et al.
@@
 Researchers have also begun to explore the use of large language models (LLMs) in medical applications.
-LLMs have been used to extract clinical concepts (Luo et al., 2022; Yang et al., 2022, 2021), and facilitate medical question answering (Yagnik et al., 2024; Singhal et al., 2023).
-However, the integration of LLMs into ED workflows, particularly in addressing ED crowding, remains unexplored.
+LLMs have been used to extract clinical concepts (Yang et al., 2021; Luo et al., 2022; Yang et al., 2022), and facilitate medical question answering (Singhal et al., 2023; Yagnik et al., 2024).
 3 MIMIC-ED-Assist Benchmark
-We discuss how we curate MIMIC-ED-Assist in collaboration with clinicians, a benchmark with comprehensive laboratory information that allows for the development of AI systems that can tackle long ED wait-times via minimizing laboratory testing time-cost.
+In this section, we discuss the curation of MIMIC-ED-Assist in collaboration with ED clinicians.
 Task Description.
-We focus on the following two prediction tasks.
-1) Flagging patients at high risk of critical outcome (i.e., death and ICU transfer), and lengthened ED.
-2) Providing time cost-effective diagnostic assistance by recommending the next most medically informative laboratory group while simultaneously minimizing the time-cost of these groups.
-As discussed in detail below, both tasks help clinicians optimize their workflow, and serve to achieve the larger goal of reducing ED LOS.
+We consider the following two tasks relevant to reducing ED LOS.
+1) Flagging patients at high risks of critical outcome (i.e., death and ICU transfer), and ED LOS exceeding $24$ hours.
+2) Providing time-cost-effective diagnostic assistance by recommending the next most medically informative laboratory group while simultaneously minimizing the time-cost of these groups.
 Data Pre-processing.
-We filter out non-hospitalized ED patients from MIMIC-IV-ED, as laboratory results are only available for admitted patients.
+Since laboratory results are only available for admitted patients, we filter out non-hospitalized ED patients from MIMIC-IV-ED.
 We only focus on adults and remove patients younger than $18$ years old.
-Further, we exclude patients with missing triage information.
+We also exclude patients that miss triage information.
 This step is necessary since clinicians order laboratory groups depending on triage information.
-We simplify our task by removing patients who receive the same test multiple times, approximately $1.5\%$ of all patients.
+To simplify the task, we remove patients who receive the same test multiple times, approximately $1.5\%$ of all patients.
 Clinically Relevant Outcomes.
-We construct two prediction outcomes that help optimize ED workflow and reduce ED LOS.
-(1) Critical condition, which refers to death during hospitalization or transfer to an ICU within 12 hours.
+In collaboration with ED clinicians, we chose the two following prediction targets.
+(1) Critical outcome, which refers to death during hospitalization or transfer to an ICU within 12 hours.
+Identifying patients at high risks of critical outcome allows clinicians to prioritize treatment and resources for them.
 (2) Lengthened ED stay, indicating if ED LOS exceeding $24$ hours.
-These outcomes were chosen as prediction targets in close collaboration with ED clinicians for the following reasons.
-If a patient is identified to be at high-risk of a critical outcome, clinicians will prioritize treatment and resources for this patient.
 Lengthened ED stay is typically correlated with the complexity of a patient’s case.
-Flagging patients at high risk of lengthened ED stay can enable timely intervention, and reduce ED LOS.
-While these two outcomes are correlated, they cover different aspects of patient care.
-For example, patients at high risk of critical outcomes often should be hospitalized quickly, while patients with complications do not necessarily have severe cases.
-As such, healthcare providers often require different diagnostics and resource management for each of these two tasks.
+Flagging patients at high risks of lengthened ED stay can enable timely intervention, and reduce ED LOS.
+The proportion of patients with these outcomes is described in Table 1.
+While these two outcomes are correlated, they also cover different aspects of patient care.
+For example, patients at high risks of critical outcomes often should be hospitalized quickly, while patients with complications do not necessarily have severe cases.
+As such, healthcare providers often require different diagnostics and resource managements for these two tasks.
 Triage Feature Selection.
-We select a number of triage features in close collaboration with ED clinicians.
-Triage features are available for every patient before any laboratory tests are ordered.
-Specifically, our ED clinicians chose $9$ triage variables that are available at the beginning of the patient encounter.
-These variables included patient demographics, medical history, vital signs, and chief complaints from triage (i.e., natural language description of symptoms).
+Triage features are measurements that are available for every patient before laboratory tests are ordered.
+We select a number of triage features in collaboration with ED clinicians.
+Specifically, we chose $9$ triage variables available at the beginning of patient encounters, which include patient demographics, medical history, vital signs, and chief complaints (i.e., natural language description of symptoms).
 Laboratory Test Selection.
 We only include laboratory tests performed in the ED.
-For simplicity, we exclude laboratory tests received by less than $5\%$ of patients and leave further examination of rare laboratory tests in future research.
+For simplicity, we exclude tests received by less than $5\%$ of patients and leave examination of rare tests to future research.
 This process results in a total of $68$ available laboratory tests.
-While there are $68$ laboratory tests, ED clinicians rarely order individual tests for a patient.
-Rather, they order groups or groups of tests based on a patient’s signs, symptoms, and risk factors.
-For example, a clinician might order a basic metabolic group or a complete blood count.
-To reflect this clinical practice, we work with ED clinicians to categorize these 68 tests into 12 distinct groups, representing common groupings in clinical settings.
-See Appendix A for all laboratory tests, and their assigned groupings.
-On average each patient has $4.7$ laboratory groups.
-Consequently, our dataset contains numerous missing values, each representing a laboratory test not administered to a specific patient.
+While there are $68$ tests, ED clinicians rarely order individual tests for a patient.
+Typically, they order groups of tests (e.g., complete blood count) based on a patient’s signs, symptoms, and risk factors.
+To reflect this clinical practice, our clinical collaborators categorized these 68 tests into 12 distinct groups.
+See Appendix A for all $68$ tests, and their assigned groupings.
+On average each patient receives $4.7$ laboratory groups.
+Consequently, MIMIC-ED-Assist contains numerous missing values, each representing a laboratory test not administered to a specific patient.
 ED LOS.
@@
 ED LOS depends on the time-costs of the administered groups and can be modeled in different ways.
-For example, sequential tests result in ED LOS being the sum of time costs, whereas parallel tests imply ED LOS is equal to the group with the largest time-cost.
+For example, sequential tests result in ED LOS being the sum of time-costs, whereas parallel tests imply ED LOS is equal to the group with the largest time-cost.
 MIMIC-ED-Assist does not specify how to approximate ED LOS from these time-costs, and instead provides this flexibility to researchers and practitioners.
 Table 1: Statistics of MIMIC-ED-Assist.
-It includes information from patient triage and lab tests.
-Lab tests are grouped into 12 groups by ED clinicians based on how they are commonly ordered.
+It includes information from patient triage and laboratory tests.
+Laboratory tests are grouped into 12 groups by ED clinicians based on how they are commonly ordered.
 An ED visit has a critical outcome if the patient is transferred to ICU or there is an inpatient mortality.
 ED stay is lengthened if the length of stay (LOS) exceeds 24 hours.
+The positive rate is shown in parentheses.
 Variable/Label Count
@@
 ##\## of triage variables 9
-##\## of lab variables 67
-##\## of lab groups 12
-Avg ##\## of lab groups per patient 4.7
+##\## of laboratory variables 67
+##\## of laboratory groups 12
+Avg. ##\## of laboratory groups per patient 4.7
 ##\## of Inpatient mortality 467 (1.44%)
@@
 Data Availability.
-Our pipeline to curate MIMIC-ED-Assist from the MIMIC-IV datasets will be made available.
-MIMIC-IV is publicly available, and then one can use our pipeline to create MIMIC-ED-Assist.
+Our pipeline to create MIMIC-ED-Assist from the MIMIC-IV dataset can be found at https://github.com/cxcscmu/ED-Copilot.
+After completing a training course and signing a data use agreement regarding patient information privacy, individuals will gain access to MIMIC-IV and can utilize our pipeline to create MIMIC-ED-Assist.
 Limitations.
-There are several limitations of MIMIC-ED-Assist, mainly due to data availability challenges in public patient data (1) MIMIC-ED-Assist is derived from MIMIC-IV which only has laboratory results for hospitalized ED patients.
+Since MIMIC-ED-Assist is derived from public patient data, it suffers from some downsides due to a lack of data availability.
+(1) MIMIC-ED-Assist is derived from MIMIC-IV which only has laboratory results for hospitalized ED patients.
 Thus, our dataset is not reflective of the entire population that visits the ED, but instead is biased towards those with more severe issues (hospitalized).
 (2) As an offline benchmark derived from past patient records, all data is retrospective.
-When training models to recommend future laboratory groups, one can only use laboratory tests patients actually received, otherwise no testing results are available.
-This leads to common challenges of offline benchmarks and may result in models learning sub-optimal laboratory group assignment, and unobserved confounding.
+As a result when developing diagnostic assistant, one can only use laboratory tests patients actually received, otherwise no testing results are available.
+This leads to common challenges of offline benchmarks and may result in models learning sub-optimal recommendations, and unobserved confounding.
+We note that measuring ‘online’ performance of ED-Copilot and AI systems in general will require clinical trials which is out of the scope of this paper.
 (3) MIMIC-IV is collected at the Beth Israel Deaconess Medical Center, and may not reflect the distribution of other healthcare systems.
-Caution is needed when generalizing models and insights derived from MIMIC-IV to other places.
+Caution is needed when generalizing models and insights derived from MIMIC-ED-Assist to new healthcare systems.
 4 ED-Copilot for Diagnostic Assistance
-With MIMIC-ED-Assist, we develop ED-Copilot, a method to provide (time)-cost-effective diagnostic assistance for ordering informative laboratory groups to flag patients at high risk for critical outcome and lengthened ED stay.
-We first provide a high-level description of ED-Copilot, followed by how it is trained.
-Training ED-Copilot involves a supervised fine-tuning stage for adapting the pre-trained language model (PLM) (e.g.
-BioGPT) backbone to our task, followed by an RL stage that selects laboratory groups to reduce time-cost.
-At the end of this section, we discuss how ED-Copilot conducts inference.
-Figure 1: Overview of ED-Copilot training flow on one ED visit.
+We use MIMIC-ED-Assist to develop ED-Copilot, which offers cost-effective diagnostic assistance by minimizing the number of laboratory groups required to identify high-risk patients.
+This section is organized as follows.
+First, we provide a high-level overview of ED-Copilot, see Figure 1 for a visualization.
+Then, we detail its training process which consists of two stages: supervised fine-tuning to adapt the pre-trained language model (PLM) (e.g., BioGPT) to our prediction task, followed by RL to select laboratory groups that reduce time-costs.
+Finally, we discuss how ED-Copilot conducts inference.
+Figure 1: Overview of ED-Copilot procedure on one ED visit.
 4.1 Problem Formulation
-Task.
-The task for ED-Copilot is to assist doctors by recommending the next laboratory group to flag patients with potential critical outcome or lengthened ED stay, based on patient triage and laboratory information.
-Let a patient have one triage group $x_{0}$ and $n$ laboratory groups $[x_{i}]_{i=1}^{n}$with their respective associated results $[r_{i}]_{i=0}^{n}$ in the order of $[x_{0},r_{0},x_{1},r_{1},x_{2},r_{2},....,x_{n},r_{n}]$, along with a diagnostic outcome $y$.
-ED-Copilot first linearizes this patient’s triage and laboratory into a text sequence, then passes it into a language model backbone to suggest the next lab group or predict the outcome.
+Let a patient has triage features denoted by $x_{0}$ and $n$ laboratory groups $[x_{i}]_{i=1}^{n}$with their associated results $[r_{i}]_{i=0}^{n}$ in order $[x_{0},r_{0},x_{1},r_{1},x_{2},r_{2},....,x_{n},r_{n}]$.
+Additionally, let $y$ denotes our prediction targets defined in Section 3.
+ED-Copilot first linearizes this patient’s triage and laboratory results into a text sequence, which is then inputted into the PLM backbone to suggest the next group of tests or predict the outcome.
+We describe these steps in detail as follows.
 […]
 Language Model Backbone.
-ED-Copilot uses a PLM $G_{\theta}$ to take the input sequence, append a [EOS] token after each group, and uses classification head upon the hidden representations of $G_{\theta}$ to select next laboratory group or predict the outcome:
+We apply PLM $G_{\theta}$ to the text sequence constructed above to obtain hidden representations H for each [EOS] token.
+We then use a MLP $p_{\phi}(x_{i}|\textbf{{h}}_{<i})$ to predict the next laboratory group using hidden representations $\textbf{{h}}_{<i}$ for tokens [EOS]<i.
+Lastly, we predict outcomes $y$ using $p_{\psi}(y|\textit{{h}}_{\leq n})$ where $\textit{{h}}_{n}$ is the hidden state for token [EOS]n.
+This process is reflected below.
 $,\displaystyle[x_{0},r_{0},\texttt{[EOS]}_{0},$$...,x_{n},r_{n},\texttt{[EOS]}_{n},y]\xrightarrow{G_{\theta}}\textit{{H}},$ (1)
@@
 $p_{\psi}(y|\textit{{h}}_{\leq n})$ $=\text{MLP}_{\psi}(\textit{{h}}_{n}).$ (3)
-Here $G_{\theta}$ generates hidden representations for each token, on which we use a multi-layer perceptron (MLP) $p_{\phi}(x_{i}|\textbf{{h}}_{<i})$ to predict the next laboratory group from the laboratory label space and another MLP $p_{\psi}(y|\textbf{{h}}_{\leq n})$ to predict the diagnostic outcome $y$, conditioned on the hidden representation h of the previous [EOS] position.
-Note that we only predict laboratory groups, not their associated laboratory tests’ results, which should be filled by running the actual lab testing.
+Note that we only predict laboratory groups, not their associated laboratory tests’ results, which should be determined by conducting the actual laboratory tests.
+Next, we describe our training procedure.
 4.2 Supervised Fine-tuning
-The first stage of training uses standard language model fine-tuning techniques, where we train the PLM to suggest the next laboratory group to order and eventually to predict target outcomes.
-To predict the next laboratory group, we use a standard auto-regressive loss function:
+In this section, we perform supervised fine-tuning of our PLM to suggest the next laboratory group and predict outcomes.
+To predict the next laboratory group, we use a standard auto-regressive loss function,
 $\mathcal{L}_{\text{lab}}$ $=-\frac{1}{n}\sum_{i=1}^{n}\log p_{\phi}(x_{i}|\textit{{h}}_{<i}).$ (4)
-For outcome prediction, we use the following loss
+For outcome prediction, we use the following loss,
 $\mathcal{L}_{y}$ $=-\log p_{\psi}(y|\textit{{h}}_{\leq n}).$ (5)
-Then, we fine-tune the PLM by:
-$\min_{\theta,\psi,\phi}(\mathcal{L}_{\text{lab}}+\mathcal{L}_{y}),$ (6)
+To fine-tune, we minimize the loss,
+$\theta^{*},\psi^{*},\phi^{*}=\min_{\theta,\psi,\phi}(\mathcal{L}_{\text{lab}}+\mathcal{L}_{y}),$ (6)
 where $\theta,\psi,\phi$ are parameters in the PLM $G_{\theta}$ and two MLPs $p_{\psi},p_{\phi}$.
@@
 4.3 Reinforcement Learning
-Choosing informative laboratory groups is a key factor in reducing laboratory testing, and hence improving wait-time, and reducing ED LOS.
-In this section, we employ RL to introduce the notion of time-cost effectiveness to the supervised fine-tuned PLM, to select laboratory groups that maximize F1-score while minimizing time-cost.
+Choosing informative laboratory groups is a key factor in reducing laboratory testing, and thereby reducing ED LOS.
+In this section, we employ RL to introduce the notion of time-cost effectiveness to the fine-tuned PLM to select laboratory groups that maximize predictive accuracy while minimizing time-cost.
 Markov Decision Process.
@@
 Action Space.
-The action space $\mathcal{A}=\{x_{1},x_{2},....,x_{K}\}\cup\{y^{+},y^{-}\}$ include $K$ number of groups, with action $a\in\{x_{1},x_{2},....,x_{K}\}$ corresponding to ordering a laboratory test group, with a time-cost $c(a)$.
-Action $a\in\{y^{+},y^{-}\}$ means predicting the outcome and terminating the process.
+Denote the action space as $\mathcal{A}=\{x_{1},x_{2},....,x_{K}\}\cup\{y^{+},y^{-}\}$.
+Action $a\in\{x_{1},x_{2},....,x_{K}\}$ corresponding to ordering a group of laboratory tests with associated time-cost $c(a)$.
+Action $a\in\{y^{+},y^{-}\}$ refers to predicting an outcome and terminating the MDP.
 Policy.
-The policy $\pi_{\eta}:\mathcal{S}\rightarrow\mathcal{A}$ maps from states to actions, and is parameterized by weights $\eta$.
+The policy $\pi_{\eta}:\mathcal{S}\rightarrow\mathcal{A}$ maps from states to actions, and is parameterized by $\eta$.
 Specifically, $\pi_{\eta}(a|\textbf{{h}}_{<i})$ outputs probability of an action $a$ using the hidden representation $\textbf{{h}}_{<i}$ corresponding to [EOS] token for state $s_{i}$.
-RL Training. We train the policy $\pi_{\eta}$ to follow two objectives: maximize F1-score and minimize time-cost. We measure the time cost of policy $\pi_{\eta}$, as follows:
+RL Training. We train the policy $\pi_{\eta}$ to follow two objectives: maximize F1-score and minimize time-cost. We measure the time-cost of policy $\pi_{\eta}$, as follows:
 $\text{Cost}(\pi_{\eta})=\mathbb{E}_{\pi_{\eta}}[\sum_{t\geq 0}\sum_{j\in[K]}c(j)\cdot\mathbf{1}\{a_{t}=x_{j}\}],$ (7)
 where $\mathbb{E}_{\pi_{\eta}}$ is the expectation under policy $\pi_{\eta}$.
-Using this definition of cost, let $\pi^{*}_{\eta}(\alpha,\beta)$ represent the policy that maximizes F1-score while minimizing time-cost for hyper-parameters $\alpha,\beta$ to be defined.
+Using (7), let $\pi^{*}_{\eta}(\alpha,\beta)$ represent the policy that maximizes F1-score while minimizing time-cost for hyper-parameters $\alpha,\beta$ to be defined.
 Then, $\pi_{\alpha,\beta}^{*}$ is equivalent to the solution of the following program:
-$\pi^{*}_{\eta}(\alpha,\beta)=\text{argmax}_{\pi_{\eta}}\{\text{TN}(\pi_{\eta})+\alpha\text{TP}(\pi_{\eta})+\beta\text{Cost}(\pi_{\eta})\}.$ (8)
-TN$(\pi_{\eta})$ and TP$(\pi_{\eta})$ are the numbers of true negatives and true positives, under policy $\pi_{\eta}$.
+$\begin{split}\pi^{*}_{\eta}(\alpha,\beta)=\text{argmax}_{\pi_{\eta}}\{&\text{TN}(\pi_{\eta})+\alpha\text{TP}(\pi_{\eta})\\ &+\beta\text{Cost}(\pi_{\eta})\}.\end{split}$ (8)
+TN$(\pi_{\eta})$ and TP$(\pi_{\eta})$ are the numbers of true negatives and true positives under policy $\pi_{\eta}$.
 Hyper-parameters $\alpha,\beta$ control the trade-off between F1-score and time-cost.
@@
 In the ED, laboratory groups are often ordered both in parallel, and sequentially since the decision on new tests to order depends on previous tests.
-Additionally, the number of tests an ED clinician orders also depends on other factors outside of the patient’s health record, such as insurance policy.
-As the first step towards reducing ED LOS with AI, this work uses the sum of test time as the optimization target, which is an approximation of the influence of test costs on total ED LOS.
+Additionally, the number of tests an ED clinician orders depends on other factors outside of the patient’s health record, such as insurance policy.
+As the first step towards reducing ED LOS with AI, we use the sum of laboratory group time-costs as the total cost, which serves as an approximation to the effect of testing on ED LOS.
 Better ways to model ED LOS are future research for both the AI and healthcare communities.
-Table 2: Overall performances of ED-Copilot and baselines in predicting Critical Outcome (transfer to ICU or mortality) and Lengthened ED Stay (spending more than 24 hours in emergency room), and the Average Time Cost (in minutes) to make a prediction.
-Sensitivity and specificity are the true positive and true negative rates.
+Table 2: Overall performance of ED-Copilot and baselines in predicting Critical Outcome (transfer to ICU or mortality) and Lengthened ED Stay (exceeding 24 hours in the ED), as well as the Average Time-cost (in minutes) to perform laboratory tests.
+Sensitivity and specificity are true positive and true negative rates.
+We report results averaged over three random seeds alongside standard deviations.
 Critical Outcome Lengthened ED Stay
-Model F1 AUC Sensitivity Specificity Avg. Time Cost F1 AUC Sensitivity Specificity Avg. Time Cost
-Random Forest 0.377 0.807 0.754 0.748 265 Minutes 0.206 0.698 0.693 0.616 265 Minutes
-XGBoost 0.379 0.807 0.731 0.744 265 Minutes 0.212 0.679 0.619 0.661 265 Minutes
-LightGBM 0.394 0.813 0.725 0.769 265 Minutes 0.217 0.705 0.706 0.605 265 Minutes
-3-layer DNN 0.339 0.743 0.676 0.683 265 Minutes 0.194 0.637 0.649 0.593 265 Minutes
-SM-DDPO 0.353 0.780 0.685 0.763 182 Minutes 0.183 0.619 0.472 0.739 177 Minutes
-ED-Copilot 0.413 0.820 0.750 0.779 125 Minutes 0.252 0.707 0.725 0.606 154 Minutes
+Model F1 AUC Sensitivity Specificity Avg. Time-cost F1 AUC Sensitivity Specificity Avg. Time-cost
+Random Forest 0.377 (0.015) 0.807 (0.011) 0.754 (0.012) 0.748 (0.005) 265 Min 0.206 (0.014) 0.698 (0.011) 0.693 (0.016) 0.616 (0.024) 265 Min
+XGBoost 0.379 (0.019) 0.807 (0.009) 0.731 (0.017) 0.744 (0.006) 265 Min 0.212 (0.010) 0.679 (0.007) 0.619 (0.020) 0.661 (0.020) 265 Min
+LightGBM 0.394 (0.016) 0.813 (0.008) 0.725 (0.012) 0.769 (0.004) 265 Min 0.217 (0.015) 0.705 (0.011) 0.706 (0.017) 0.605 (0.014) 265 Min
+3-layer DNN 0.339 (0.032) 0.743 (0.021) 0.676 (0.024) 0.683 (0.011) 265 Min 0.194 (0.031) 0.637 (0.013) 0.649 (0.015) 0.593 (0.014) 265 Min
+SM-DDPO 0.353 (0.031) 0.780 (0.020) 0.685 (0.023) 0.763 (0.022) 182 (32) Min 0.183 (0.028) 0.619 (0.012) 0.472 (0.012) 0.739 (0.011) 177 (60) Min
+ED-Copilot 0.413 (0.028) 0.820 (0.021) 0.750 (0.018) 0.779 (0.011) 125 (21) Min 0.232 (0.023) 0.707 (0.015) 0.725 (0.018) 0.606 (0.015) 154 (33) Min
 4.4 Inference
-During inference, ED-Copilot assists ED clinicians to optimize their workflow by suggesting the next most informative laboratory group to order, and also by flagging patients at high risk of critical outcome and lengthened ED stay.
+During inference, ED-Copilot assists clinicians to optimize their workflow by suggesting the next most informative laboratory group, and also by flagging patients at high risks of critical outcome and lengthened ED stay.
 Specifically, given triage information and previous test results, ED-Copilot recommends additional tests.
-The results for these additional tests can then be fed back to ED-Copilot to suggest additional laboratory groups or to flag one of two possible outcomes.
-Further, ED-Copilot can be used in more flexible ways depending on a clinician’s needs.
+The results of these tests are fed back to ED-Copilot to suggest additional laboratory groups or to flag one of two possible outcomes.
+Further, ED-Copilot can be used in more others ways depending on clinical needs.
 For example, multiple tests can be ordered using multiple suggestions from ED-Copilot.
@@
 (d) Lengthened ED Stay AUC
-Figure 2: Prediction accuracy and average number of laboratory groups of ED-Copilot with different maximum allowed times to perform lab tests.
-Each point reflects ED-Copilot’s F1/AUC (y-axes) at different time thresholds.
+Figure 2: Prediction accuracy and average number of laboratory groups of ED-Copilot with different maximum allowed time to perform laboratory tests.
+Each point reflects ED-Copilot’s F1/AUC (y-axes) at different time upper-bounds.
 5 Experimental Set-up
@@
 We randomly split the dataset using $80\%$ for training, $10\%$ for validation, and $10\%$ for testing, while ensuring each split has the same class distribution.
-The validation set is used to tune hyper-parameters for all methods.
-During inference, the initial state of that patient is set to be demographic and triage information.
+The validation set is used to tune hyper-parameters.
+During inference, the initial state of a patient is set to their triage information.
 Evaluation Metrics.
-We use four evaluation metrics: F1-score, area under the receiver operating characteristic (AUC), sensitivity (true positive rate), and specificity (true negative rate), all standard in healthcare tasks prediction tasks (Harutyunyan et al., 2019; Xie et al., 2022).
+We use four evaluation metrics: F1-score, area under the receiver operating characteristic (AUC), sensitivity (true positive rate), and specificity (true negative rate) which are all standard in healthcare tasks prediction tasks (Harutyunyan et al., 2019; Xie et al., 2022).
 Baselines.
 For prediction tasks only, ED-Copilot is compared to three tree-based methods: random forests (Breiman, 2001), XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017), all of which are known to achieve strong performance on tabular data.
-Additionally, we compare ED-Copilot to a $3$-layer dense deep neural network (DNN).
-Note that since each patient is not observed under all lab groups, there are many missing values in our dataset.
-For these baselines, we consider a number of imputation methods: (1) mean imputation, (2) median imputation, and (3) zero-imputation (i.e., replacing missing values by 00), (4) using a dummy indicator to encode missing values.
-Results with the best imputation method is presented in the main manuscript, and the rest are in Appendix C.
-We also compare ED-Copilot to another cost-effective baseline, SM-DDPO (Yu et al., 2023), which selects laboratory groups to minimize time cost while maximizing F1-score.
-SM-DDPO is a non-personalized method, and selects the same laboratory groups for all patients.
-All hyper-parameters are tuned on the validation set to maximize F1-score.
-Time Cost.
-Each group of lab tests is assigned a time cost by observing time-stamps in the MIMIC-IV database.
-For both ED-Copilot, and all baselines, we measure the amount of time taken by averaging the time required for an algorithm to make a prediction across all patients.
-This reflects the impact of laboratory tests on ED LOS.
+We also compare ED-Copilot to a $3$-layer dense deep neural network (DNN).
+Since each patient is not observed under all laboratory groups, there are many missing values in MIMIC-ED-Assist.
+Since baselines cannot naturally handle missing values, we consider a variety of imputation methods: (1) mean imputation, (2) median imputation, and (3) zero-imputation (i.e., replacing missing values by 00), (4) using a dummy indicator to encode missing values.
+Results for the best imputation method are presented in the main manuscript.
+The rest are in Appendix C.
+We compare ED-Copilot to another cost-effective baseline, SM-DDPO (Yu et al., 2023), which selects laboratory groups to minimize time-cost while maximizing F1-score.
+SM-DDPO is a non-personalized method, i.e., it selects the same laboratory groups for all patients.
+Time-cost.
+As discussed, each group of tests is assigned a time-cost by observing time-stamps in the MIMIC-IV database.
+For all methods, we measure the amount of time taken by averaging the time required for an algorithm to make a prediction across patients.
 Implementation Details.
-BioGPT (Luo et al., 2022) is used as our backbone language model, which is a generative pre-trained transformer (347 million parameters) for biomedical text generation and mining.
+BioGPT (Luo et al., 2022) is used as our backbone language model, which is a generative pre-trained transformer (347 million parameters) for biomedical text generation and representation.
 In the RL phase of training ED-Copilot, for a given patient, our action space is restricted to laboratory groups that the patient actually performed.
-A discussion when selecting unobserved laboratory groups can be found in Section 6.4.
+Experiments when ED-Copilot is allowed to select unobserved laboratory groups can be found in Section 6.5.
 Table 3: Ablation study on components of ED-Copilot, including linearization techniques, feature importance, and PLM backbones.
 Group Variations Critical Outcome Lengthened ED Stay
-F1 AUC Sensitivity Specificity Avg. Time Cost F1 AUC Sensitivity Specificity Avg. Time Cost
-ED-Copilot (345M) 0.413 0.820 0.750 0.779 125 Minutes 0.252 0.707 0.725 0.606 154 Minutes
-Linearization Raw Lab Test Name 0.397 0.777 0.768 0.677 134 Minutes 0.241 0.695 0.611 0.701 144 Minutes
+F1 AUC Sensitivity Specificity Avg. Time-cost F1 AUC Sensitivity Specificity Avg. Time-cost
+ED-Copilot (345M) 0.413 0.820 0.750 0.779 125 Min 0.252 0.707 0.725 0.606 154 Min
+Linearization Raw Lab Test Name 0.397 0.777 0.768 0.677 134 Min 0.241 0.695 0.611 0.701 144 Min
 Features w/o. Triage 0.277 0.704 0.679 0.649 0.145 0.593 0.532 0.606
 w/o. CBC 0.385 0.803 0.692 0.777 0.224 0.686 0.696 0.596
-w/o. Chemical 0.420 0.827 0.788 0.746 0.234 0.702 0.656 0.606
-Backbone Pythia (70M) 0.290 0.698 0.574 0.702 166 Minutes 0.178 0.596 0.555 0.619 126 Minutes
-GPT2-Medium (345M) 0.358 0.757 0.621 0.747 133 Minutes 0.166 0.539 0.498 0.584 96 Minutes
+w/o. CHEM 0.420 0.827 0.788 0.746 0.234 0.702 0.656 0.606
+Backbone BioGPT (345M) w/o. RL 0.381 0.810 0.725 0.765 265 Min 0.236 0.718 0.710 0.620 265 Min
+Llama (7B LORA) w/o. RL 0.397 0.798 0.692 0.767 265 Min 0.232 0.701 0.705 0.610 265 Min
+[…]
+GPT2-Medium (345M) w. RL 0.358 0.757 0.621 0.747 133 Min 0.166 0.539 0.498 0.584 96 Min
 6 Evaluation Results
-6.1 Prediction Accuracy and Time Cost
+In this section, we present our experimental results.
+Section 6.1 compares the predictive and time-cost performance of ED-Copilot with baselines.
+Section 6.2 describes our ablation studies.
+Sections 6.3 and 6.4 establish ED-Copilot’s ability to personalize recommendations based on patient severity and achieve consistent performance across subgroups such as age and sex.
+Lastly, in Section 6.5, we investigate ED-Copilot’s performance when it is not restricted to select administered groups.
+6.1 Prediction Accuracy and Time-cost
 Overall Performance.
-The prediction performance and time costs of all methods are presented in Table 2.
+The prediction performance and time-costs of all methods are presented in Table 2.
 For critical outcome, ED-Copilot outperforms the next best baseline by 1.9%, 0.7%, 2.5%, and 1% for F1-score, AUC, sensitivity, and specificity, respectively.
 For lengthened ED stay, ED-Copilot outperforms the next best model by 3.5%, 0.2%, 1.9%, and 0.1% for the four metrics.
-ED-Copilot halves average time cost from roughly $4$ hours to $2$ hours.
-In comparison to the other cost-effective baseline, SM-DDPO, ED-Copilot also has significantly better accuracy and lower time cost.
-Both methods utilize the same RL algorithm to minimize time-cost, but ED-Copilot benefits from the strong capability of PLM backbone (Sec 6.2) and the personalized assistance it enabled (Sec 6.3).
-Prediction Accuracy at Different Time Costs.
+ED-Copilot halves average time-costs from roughly $4$ hours to $2$ hours.
+In comparison to the other cost-effective baseline, SM-DDPO, ED-Copilot also has significantly better accuracy and lower time-costs.
+Both methods utilize the same RL algorithm to minimize time-cost, but ED-Copilot benefits from the strong capability of PLM backbone (Sec 6.2) and the personalized diagnostic assistance it enabled (Sec 6.3).
+Prediction Accuracy at Different Time-costs.
 In Figure 2, we investigate how ED-Copilot’s prediction performance changes as we vary the maximum time allowed to perform laboratory tests.
-The prediction performance increases with the maximum allowed time and number of laboratory groups selected.
-ED-Copilot requires relatively few tests to approximate peak performance for lengthened ED stay, while prediction accuracy for critical outcome steadily increases with more laboratory tests allowed.
-These curves illustrate different trade-offs of diagnostic assistance for different outcomes and help ED clinicians better leverage when allocating resources.
+Performance increases with maximum allowed time and number of laboratory groups selected.
+ED-Copilot requires relatively few tests to achieve peak performance for lengthened ED stay, while prediction accuracy for critical outcome steadily increases as more tests are allowed.
+These curves illustrate different requirements for different outcomes and shed light as to how clinicians should allocate resources across tasks.
+6.2 Ablation Studies
+Linearization Technique.
+Following Hegselmann et al.
+(2023) and Manikandan et al.
+(2023), we replace the true laboratory test name with a standard feature name while keeping the result for that laboratory group unchanged.
+That is, we linearize laboratory results as follows ( feature1 : value1 |||| feature2 : value2 … ).
+Table 3 shows using standard feature names leads to better F1 and AUC score as compared to using the raw laboratory test names.
+One possible reason is that using raw laboratory test names biases ED-Copilot to pick names over informativeness for prediction.
+Feature Importance.
+We examine the change in ED-Copilot performance when training with and without a given feature set from top three popular feature sets: triage information, complete blood count (CBC), and chemical group (CHEM).
+Table 3 shows removing triage features leads to a significant drop in performance.
+Our ED collaborators confirms this is medically plausible since initial laboratory tests are based on triage information.
+Our results also show all laboratory groups are not equally important, for example, removing CBC causes a larger drop in performance than removing the chemical laboratory group.
 (a) Sensitivity-Specificity $\alpha$
@@
 Figure 3: Impact of Hyper-parameters on Sensitivity-Specificity ($\alpha$) and F1-Cost ($\beta$) trade-off when predicting critical outcome.
-6.2 Ablation Studies
 PLM Backbone.
-We compare BioGPT to GPT2 at the same size, and a smaller scale model Pythia (Biderman et al., 2023) that has 70 million parameters.
+We compare BioGPT to GPT2 which is the same size, and a smaller model Pythia (Biderman et al., 2023) that has 70 million parameters.
 BioGPT is based on GPT2 (Radford et al., 2019), but is further pre-trained on a bio-medical corpus.
 Table 3 shows both model scale and bio-medical pre-training lead to stronger performance.
-Linearization Technique.
-As done in Hegselmann et al.
-(2023); Manikandan et al.
-(2023), we replace the true laboratory test name with a standard feature name while keeping the result for that laboratory group unchanged.
-That is, we linearize laboratory results as follows ( feature1 : value1 |normal-||| feature2 : value2 … ).
-Table 3 shows using standard feature names leads to better F1 and AUC score as compared to using the raw laboratory test names.
-One possible reason is that using raw laboratory test names biases ED-Copilot to pick names over informativeness for prediction.
-Feature Importance.
-We examine the change in ED-Copilot performance when training with and without a given feature set from top three popular feature sets: triage information, complete blood count, and chemical group.
-Table 3 shows removing triage features leads to a significant drop in performance.
-Our ED collaborators confirmed this made medical sense since initial laboratory tests are based on triage information.
-Our results also show all laboratory groups are not equally important, with removing the CBC causing a larger drop in performance than excluding the chemical laboratory groups.
-Figure 4: Fraction of patients performing lab groups and predicted by ED-Copilot. On average each patient performed 4.7 groups and cost-effective ED-Copilot suggested 2.4 groups.
+We also compare Llama (7B) (Touvron et al., 2023) without RL for this task to BioGPT (345M) without RL, and ED-Copilot.
+Llama is fine-tuned using Low-Rank Adaptation (LORA) (Hu et al., 2022) for efficiency.
+Llama 7B (LORA) without RL performs comparably to BioGPT without RL.
+This indicates that scaling up the backbone parameters is promising avenue to boost absolute performance.
+However, Llama 7B (LORA) underperforms BioGPT with RL (i.e., ED-Copilot), particularly when it comes to time-cost.
+This is expected as the PLM backbone and the RL serve different roles in the system.
+PLM is used to predict outcomes while RL minimizes time-cost.
+Fine-tuning does not address time-costs.
+These two aspects are both necessary.
 Hyper-parameters.
@@
 Figure 3(a) shows increasing $\alpha$ leads to trade off sensitivity over specificity.
-Figure 2(b) show increasing $\beta$ causes ED-Copilot to trade off F1-score over time-cost.
+Figure 3(b) shows increasing $\beta$ causes ED-Copilot to trade off F1-score over time-cost.
 The hyper-parameters $(\alpha,\beta)$ provide a control for ED clinicians to change ED-Copilot’s behavior to fit their preferences.
-6.3 Analysis on Personalized Diagnostic Assistance
-Table 4: Critical Outcome prediction accuracy at different cohorts, by the category of lab groups they performed (Top: Performed at least one of the top two most frequent lab groups; Middle: Performed at least one of the middle six lab groups; Rare: Performed one or more rare tests in the last four lab groups).
+Figure 4: Fraction of patients performing laboratory groups and predicted by ED-Copilot. On average each patient performed 4.7 groups and cost-effective ED-Copilot suggested 2.4 groups.
+Table 4: Critical Outcome prediction accuracy at different cohorts, by the category of laboratory groups they performed (Top: Performed at least one of the top two most frequent laboratory groups; Middle: Performed at least one of the middle six laboratory groups; Rare: Performed one or more rare tests in the last four laboratory groups).
 The total number of positive (critical)/negative cases and positive rate is shown in parentheses.
@@
 ED-Copilot 0.414 0.701 0.788 0.431 0.731 0.774 0.461 0.767 0.720
+Table 5: Model performance using only triage information under supervised fine-tuning training.
+Model Critical Outcome Lengthened ED Stay
+F1 AUC Sensitivity Specificity F1 AUC Sensitivity Specificity
+Random Forest 0.346 0.767 0.689 0.731 0.181 0.617 0.597 0.571
+XGBoost 0.324 0.736 0.673 0.668 0.165 0.603 0.580 0.560
+LightGBM 0.341 0.743 0.624 0.738 0.171 0.615 0.554 0.628
+3-layer DNN 0.275 0.712 0.611 0.702 0.155 0.553 0.485 0.597
+ED-Copilot (SFT) 0.392 0.792 0.708 0.762 0.199 0.675 0.679 0.557
+6.3 Analysis on Personalized Diagnostic Assistance
 Personalized treatment strategies are often required since patients have diverse needs and present differently.
 ED-Copilot learns a personalized representation for each patient and naturally supports personalized diagnostic assistance.
-On the other hand, previous cost-effective methods based on traditional ML algorithms (e.g., tree-based methods), optimize cost-effectiveness at the full population level, i.e., via feature selection/pruning, rather than personalized for each patient.
-However, this does not naturally support personalization, unlike ED-Copilot which suggests individualized future tests based on a patient’s triage and previous test results.
+Previous cost-effective methods based on traditional ML algorithms such as tree-based methods optimize cost-effectiveness at the full population level, e.g., via feature selection.
+Unlike ED-Copilot, this does not lead to personalized recommendations that are based on triage and previous test results.
 The following experiments highlight the benefit of personalization via ED-Copilot.
-First, we plot both the fraction of patients receiving each group as well as the fraction of patients predicted by ED-Copilot in Figure 4.
-In the ED, patients undergo a large variety of laboratory groups: after the two most common groups (Chem and CBC), more than half of the patients performed some other tests.
-ED-Copilot predicts fewer laboratory groups than the observed data since it is trying to minimize time-cost, yet it still predicts a wide variation of recommended tests depending on the patient’s condition.
-On the other hand, non-personalized cost-effective methods, often withdraw to picking the most frequent groups.
-To further understand the benefit of personalization, we partition patients in the test set into three cohorts based on the rarity of the group they were administered; see Table 4 for a precise definition of these cohorts.
-The middle and rare cohorts have more positive cases, indicating higher severity in these cohorts.
+We plot both the fraction of patients receiving each group of tests as well as the fraction of patients predicted by ED-Copilot to undergo each of these groups in Figure 4.
+This plot shows significant variation in tests received by ED patients.
+After the two most common groups (CHEM and CBC), more than half of the patients performed some other tests.
+ED-Copilot reflects this diversity by recommending a variety of tests based on patients’ condition.
+On the other hand, non-personalized cost-effective methods often withdraw to picking the most frequent groups.
+To further understand the benefit of personalized diagnostic assistance, we partition patients in the test set into three cohorts based on the rarity of laboratory groups they were administered.
+A precise definition of these cohorts can be found in Table 4.
+The middle and rare cohorts have a larger proportion of positive cases, indicating higher severity.
 We train tree-based baselines using triage information and results from the two most frequent groups (CHEM and CBC), and compare it to ED-Copilot.
-Results in Table 4 show that ED-Copilot and baselines perform comparably on patients in the top cohorts (i.e., patients that present typically).
-For the middle cohort, ED-Copilot greatly improves F1-score as well.
-Further, in the rare cohort which has the highest fraction of critical cases as compared to other groups (15% versus 10%), ED-Copilot achieves higher significantly higher sensitivity than other methods.
-Hence, ED-Copilot provides personalized assistance to provide more equitable care, particularly for high-risk patients.
+Results in Table 6 show ED-Copilot and baselines perform comparably on patients in the low severity cohorts (i.e., patients that present typically).
+For the middle cohort, ED-Copilot greatly improves F1-score.
+In the rare cohort which has the highest proportion of critical outcomes (15% versus 10% in middle cohort), ED-Copilot achieves higher significantly higher sensitivity than other methods.
+This indicates ED-Copilot’s ability to provide personalized assistance which results in more equitable care, particularly for high-risk patients.
+Table 6: Sub-group Analysis for Critical Outcome and Lengthened ED Stay.
+Model Critical Outcome Lengthened ED Stay
+F1 AUC Sensitivity Specificity Avg. Time-cost F1 AUC Sensitivity Specificity Avg. Time-cost
+Group: Male, Positive/Negative (Ratio): 155/1329 (10.44%) on critical outcome, 104/1380 (7.01%) on lengthened ED stay
+Random Forest 0.387 0.793 0.785 0.692 265 0.207 0.723 0.644 0.704 265
+XGBoost 0.379 0.804 0.726 0.732 265 0.258 0.722 0.692 0.659 265
+LightGBM 0.391 0.807 0.762 0.742 265 0.223 0.724 0.769 0.597 265
+ED-Copilot 0.377 0.780 0.671 0.775 121 0.242 0.670 0.574 0.702 151
+Group: Female, Positive/Negative (Ratio): 143/1608 (8.17%) on critical outcome, 127/1624 (7.25%) on lengthened ED stay
+Random Forest 0.385 0.819 0.730 0.794 265 0.193 0.677 0.646 0.616 265
+XGBoost 0.368 0.826 0.766 0.740 265 0.171 0.645 0.661 0.576 265
+LightGBM 0.394 0.813 0.766 0.730 265 0.198 0.683 0.669 0.636 265
+ED-Copilot 0.413 0.823 0.727 0.819 129 0.215 0.694 0.713 0.621 160
+Group: Age 18-30, Positive/Negative (Ratio): 22/341 (6.06%) on critical outcome, 23/340 (6.34%) on lengthened ED stay
+Random Forest 0.277 0.871 0.889 0.838 265 0.197 0.742 0.739 0.738 265
+XGBoost 0.413 0.908 0.833 0.884 265 0.306 0.764 0.652 0.824 265
+LightGBM 0.390 0.895 0.889 0.913 265 0.259 0.767 0.783 0.629 265
+ED-Copilot 0.208 0.849 0.864 0.708 106 0.304 0.823 0.956 0.705 159
+Group: Age 31-60, Positive/Negative (Ratio): 97/1169 (7.66%) on critical outcome, 92/1174 (7.27%) on lengthened ED stay
+Random Forest 0.367 0.824 0.739 0.794 265 0.196 0.675 0.620 0.670 265
+XGBoost 0.367 0.835 0.750 0.775 265 0.228 0.685 0.674 0.658 265
+LightGBM 0.400 0.835 0.802 0.756 265 0.213 0.692 0.739 0.611 265
+ED-Copilot 0.409 0.800 0.711 0.823 124 0.209 0.697 0.612 0.664 153
+Group: Age 61-90, Positive/Negative (Ratio): 189/1261 (13.03%) on critical outcome, 108/1342 (7.45%) on lengthened ED stay
+Random Forest 0.411 0.780 0.710 0.753 265 0.208 0.707 0.685 0.621 265
+XGBoost 0.383 0.785 0.659 0.789 265 0.188 0.655 0.704 0.552 265
+LightGBM 0.405 0.779 0.688 0.757 265 0.192 0.698 0.667 0.622 265
+ED-Copilot 0.400 0.749 0.667 0.694 129 0.222 0.695 0.731 0.671 156
+6.4 Sub-group Analysis
+We compare the performance of ED-Copilot to tree-based baselines across sub-groups including gender and age to study if the algorithm is working fairly across all sub-groups of interest in Table 6.
+ED-Copilot displays consistent performance, and outperforms baselines across subgroups on both prediction tasks.
+6.5 Performance of Unrestricted Lab Group Suggestion
+Since MIMIC-ED-Assist is an offline retrospective benchmark, we restrict ED-Copilot during training to only select laboratory groups that patients have received.
+However, this restriction may lead to sub-optimal group suggestion.
+We investigate ED-Copilot’s performance without this restriction through the following experiment.
 (a) Accuracy
 (b) Avg. # of Labs
-Figure 5: Comparison of prediction performance on critical outcome of ED-Copilot when restricted or not restricted to tests patients performed in its suggestion at different time cost constraint.
-6.4 Unrestricted Lab Group Suggestion
-Since MIMIC-ED-Assist is an offline retrospective benchmark, we restrict ED-Copilot during training to only select laboratory groups that patients have received.
-However, this restriction may not lead to the optimal group suggestion.
-This experiment investigates ED-Copilot’s performance without this restriction.
-Specifically, if ED-Copilot (unrestricted) selects a laboratory group not received by a patient, we impute those results by 00, but add to the time-cost.
-ED-Copilot (unrestricted) achieves an accuracy of $77\%$ when predicting the next laboratory group, i.e., the group actually received by a given patient, indicating that ED-Copilot (unrestricted) is reasonably good at fitting the clinician’s decision but there are still differences between model suggestions and historical data.
+Figure 5: Comparison of prediction performance on critical outcome of ED-Copilot when restricted or not restricted to tests patients performed in its suggestion at different time-cost constraints.
+If ED-Copilot (unrestricted) selects a laboratory group not received by a patient, we impute those results by 00, but add to the time-cost.
+ED-Copilot (unrestricted) achieves an accuracy of $77\%$ when predicting the next laboratory group, i.e., the group actually received by a given patient, indicating that ED-Copilot (unrestricted) is reasonably good at replicating clinical decisions but there are still differences between model suggestions and historical data.
 Next, we compare ED-Copilot (unrestricted) to ED-Copilot (restricted) in terms of F1-score for critical outcome when varying the maximum allowed time.
-Figure 5 show that for low maximum allowed time, ED-Copilot (unrestricted) performs worse than ED-Copilot since if it picks unobserved labs, the cost brings no new information added (no results available).
-As the number of allowed laboratory groups increases, ED-Copilot (unrestricted) matches the performance of ED-Copilot.
-Restriction to observed laboratory groups might be necessary for offline evaluation.
-When used in real scenarios, ED-Copilot would serve as a “Co-Pilot” and suggest laboratory groups from the entire set without restriction, assisting ED clinicians to optimize their workflows.
+Figure 5 shows that when maximum allowed time is small, ED-Copilot (unrestricted) performs worse than ED-Copilot.
+This is because if ED-Copilot (unrestricted) picks a laboratory test that was not administered to a patient, this selection does not result in new information being added.
+However, as the number of allowed laboratory groups increases, ED-Copilot (unrestricted) matches the performance of ED-Copilot.
+Restriction to observed laboratory groups is necessary for offline evaluation.
+In real-world settings, ED-Copilot should serve as a “Co-Pilot” and suggest laboratory groups from the entire set without restriction, and assist ED clinicians in optimizing their workflows.
+We also evaluate ED-Copilot on new patients only using triage data, and without any laboratory results which are the information available on patient arrival.
+We use ED-Copilot only with supervised fine-tuning (SFT), and without RL.
+Results are displayed in Table 5.
+Without laboratory results, all models suffer, but ED-Copilot (SFT) outperforms baselines, showing its utility at an early treatment stage.
 7 Conclusion
-In this paper, we aim at reducing emergency department wait time through time-cost-effective diagnostic assistance, in collaboration with ED clinicians.
-Specifically, we curate MIMIC-ED-Assist, a benchmark that contains comprehensive laboratory information annotated with medically relevant groupings alongside key clinical outcomes that serve as useful approximations of ED wait-time.
-Using MIMIC-ED-Assist, we propose ED-Copilot, an AI system that provides time-cost-effective diagnostic assistance for ED clinicians, by recommending informative laboratory groups, and flagging patients at high risk of critical outcome and lengthened ED stay.
-ED-Copilot outperforms baselines by significantly improving prediction accuracy and reducing laboratory time costs.
-Further, its language model backbone allows personalized diagnostic assistance to better address the needs of patients who present atypically.
-We believe that this work serves as a stepping stone towards AI-driven solutions in the ED, and we hope that MIMIC-ED-Assist will spur interest in applying advancements in ML and language models to tackle this critical healthcare challenge.
-8 Acknowledgments
+In this paper, we aim at reducing emergency department wait time through time-cost-effective diagnostic assistance.
+In collaboration with ED clinicians, we curate MIMIC-ED-Assist, a benchmark that contains comprehensive laboratory information annotated with medically relevant groupings alongside key clinical outcomes that serve as useful approximations to ED wait-time.
+Using MIMIC-ED-Assist, we propose ED-Copilot, an AI system that provides time-cost-effective diagnostic assistance for ED clinicians, recommending informative laboratory groups and flagging patients at high risks of critical outcome and lengthened ED stay.
+ED-Copilot outperforms baselines by significantly improved prediction accuracy and reduced laboratory time-costs.
+Its language model backbone allows for personalized diagnostic assistance to better address the needs of severe patients.
+We believe this work takes a step towards AI-driven assistance in the ED and hope that MIMIC-ED-Assist spurs interest in applying advancements in AI to tackle this critical healthcare challenge.
+Acknowledgments
 We gratefully acknowledge partial support from NSF grants DMS-2209975, 2015341, NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI), the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and 814639, and NIH grant R01GM152718.
 We also acknowledge support by the Eunice Kennedy Shriver National Institute of Child Health and Human Development of the National Institutes of Health under Award Number K23HD110716.
-We also thank Yikuan Li from Northwestern University to share their preprocessed sepsis dataset.
-Appendix A Triage and Lab Test Group
-In Table 5, we include triage assessments and 12 lab groups with 68 tests and their estimated time cost.
+We also thank Yikuan Li from Northwestern University to share their pre-processed sepsis dataset.
+Impact Statement
+The development of ED-Copilot marks an advancement in medical technology, offering an AI-driven diagnostic tool to improve patient care in ED.
+By expediting diagnoses, ED-Copilot has the potential to increase efficiency of ED operations and enhance patient care.
+From an ethical perspective, the deployment of ED-Copilot carries a significant responsibility on the privacy and security of patient data.
+As the system will handle sensitive health information, strict measures, for example, internal-accessible only systems and compliant data protection procedures, must be in place to protect against breaches and misuse, ensuring patient confidentiality.
+In the broader societal context, the implementation of ED-Copilot aims to address the issue of ED congestion.
+It is vital to ensure that the benefits of such technology are accessible to all segments of society, regardless of socioeconomic status.
+Equity in healthcare technology means that all patients, irrespective of their background, should have the opportunity to benefit from advancements like ED-Copilot.
+Moreover, the development of ED-Copilot should support healthcare professionals, not replace them, and should be viewed as a tool to assist medical staff, allowing them to focus on the more nuanced aspects of patient care.
+Appendix A Triage and Laboratory Test Group
+In Table 7, we include triage assessments and 12 laboratory groups with 68 tests and their estimated time-costs.
 Table 7: Triage and Lab Test Group
@@
 Toxicology (TOX) (70 min) Ethanol
-Inflammation (178 min) Creatine Kinase (CK) C-Reactive Protein
+Inflammation (INFLAM) (178 min) Creatine Kinase (CK) C-Reactive Protein
 Appendix B Training Details
@@
 Due to large class imbalance, we use class weights when training the diagnostic predictor $p_{\psi}$.
-All experiments, training and hyper-parameter tuning are conducted on a NVIDIA RTX A6000 GPU.
+All experiments, training and hyper-parameter tuning are conducted on one NVIDIA RTX A6000 GPU.
 During the RL phase of training ED-Copilot, we restrict the action space to be on laboratory groups that were administered to patients.
 That is, we do not select laboratory groups that patients have not received.
-Performance when this restriction is removed is explored in Section 6.3.
-To ensure ensure sufficient policy sampling on experience replay buffer and obtain a large batch size, we freeze the weights of our PLM, to train the policy $\pi_{\eta}$, which is parameterized by $\eta$.
-We use the PPO algorithm (Schulman et al., 2017) to train $\pi_{\eta}$.
-The loss function employed is as the following for hyper-parameters $c_{1}$, and $c_{2}$:
+To ensure sufficient policy sampling on experience replay buffer and obtain a large batch size, we freeze the weights of our PLM.
+To train the policy $\pi_{\eta}$, we use the PPO algorithm (Schulman et al., 2017).
+The loss function employed is as the following for hyper-parameters $c_{1}$ and $c_{2}$:
 $\mathcal{L}_{\text{rl}}(\eta;\eta^{old})$ $=\mathcal{L}_{\text{clip}}(\eta;\eta^{old})-c_{1}\mathcal{L}_{\text{value}}(\eta)+c_{2}\text{Entropy}[\pi_{\eta}],$
 $\mathcal{L}_{\text{clip}}(\eta;\eta^{old})$ $={\hat{\mathbb{E}}}_{i}\left[\min\left(\frac{\pi_{\eta}(a_{i}|{\textbf{{h}}}_{<i})}{\pi_{\eta^{old}}(a_{i}|{\textbf{{h}}}_{<i})}\hat{A}_{i},\text{clip}\left(\frac{\pi_{\eta}(a_{i}|{\textbf{{h}}}_{<i})}{\pi_{\eta^{old}}(a_{i}|{\textbf{{h}}}_{<i})},1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right)\right],$
-$\mathcal{L}_{\text{value}}(\eta)$ $=\hat{\mathbb{E}}_{i}\left[(V_{\eta}({\textbf{{h}}}_{<i})-V_{i})^{2}\right],$
-Here ${\textbf{{h}}}_{<i}$ is the hidden representation from PLM of last [EOS] in this patient’s linearized sequence in state $s_{i}$.
-$\mathcal{L}_{\text{clip}}$ is clipped surrogate loss and $\hat{A}_{i}$ is estimated advantages which is regularized by value function $\mathcal{L}_{\text{value}}$.
-$\text{Entropy}[\pi_{\eta}]$ denotes an entropy term over the states, and $\hat{\mathbb{E}}_{i}$ is the empirical average over the collected dataset.
-We used a masked actor-critic network with package stable-baseline3 (Raffin et al., 2021).
-See algorithm 1 for a high-level description of this method.
+$\mathcal{L}_{\text{value}}(\eta)$ $=\hat{\mathbb{E}}_{i}\left[(V_{\eta}({\textbf{{h}}}_{<i})-V_{i})^{2}\right].$
+[…]
 Algorithm 1 Proximal Policy Optimization (PPO)
@@
   end for
-We list the hyper-parameters in Table 6, including the supervised fine-tuning and reinforcement learning stage.
-In the RL stage, we use grid-search to tune $\alpha$ and $\beta$ to balance the trade-off between accuracy and cost.
-The search scope for $\alpha\in\{\frac{1}{8},\frac{1}{4},\frac{1}{2},1,2,4,8,15,16,32,64,256\}$ and $\beta\in\{\frac{1}{100},\frac{1}{50},\frac{1}{20},\frac{1}{10},1,10,100,1000\}$.
-The parameters in PPO that are not specified are assigned the default values found in the Python package (Schulman et al., 2017).
 Table 8: Hyperparameters Configurations for ED-Copilot
@@
 Epochs 15 10
-Max sequence length 768 656
+Max sequence length 656 656
 Class Weight 10 -
@@
 The penalty ratio between false positive and false negative $\alpha$ 15
-The penalty ratio between false positive and lab cost $\beta$ $\frac{1}{100}$
-Appendix C Baseline Details
-In Table 7, we present tree models’ results with different imputation methods: (1) mean imputation, (2) median imputation, and (3) zero-imputation (i.e., replacing missing values by 0), (4) using a dummy indicator to encode missing values.
+The penalty ratio between false positive and laboratory cost $\beta$ $\frac{1}{100}$
+We list the hyper-parameters in Table 8, including the supervised fine-tuning and reinforcement learning stage.
+In the RL stage, we use grid-search to tune $\alpha$ and $\beta$ to balance the trade-off between accuracy and cost.
+The search scope for $\alpha\in\{\frac{1}{8},\frac{1}{4},\frac{1}{2},1,2,4,8,15,16,32,64,256\}$ and $\beta\in\{\frac{1}{100},\frac{1}{50},\frac{1}{20},\frac{1}{10},1,10,100,1000\}$.
+The parameters in PPO that are not specified are assigned the default values found in the Python package (Schulman et al., 2017).
+Appendix C More Baseline Results
+In Table 9, we present tree models’ results with different imputation methods: (1) Mean imputation.
+(i.e.
+replacing missing values with the mean value of the non-missing data for the particular feature).
+(2) Median imputation.
+(i.e.
+replacing missing values with the median value of the non-missing data for the particular feature) (3) Zero imputation.
+(i.e.
+replacing missing values by 0).
+(4) A dummy indicator to encode missing values, which is used in XGBoost and LightGBM.
 Table 9: Results with Different Imputation Methods